### v1

# Enhancing Accuracy in Differentially Private Distributed Optimization Through Sensitivity Reduction

Abstract
In this paper, we investigate the problem of differentially private distributed optimization. Recognizing that lower sensitivity leads to higher accuracy, we analyze the key factors influencing the sensitivity of differentially private distributed algorithms. Building on these insights, we propose a novel differentially private distributed algorithm that enhances optimization accuracy by reducing sensitivity. To ensure practical applicability, we derive a closed-form expression for the noise parameter as a function of the privacy budget. Furthermore, we rigorously prove that the proposed algorithm can achieve arbitrarily rigorous $\epsilon$-differential privacy, establish its convergence in the mean square sense, and provide an upper bound on its optimization accuracy. Finally, extensive comparisons with various privacy-preserving methods validate the effectiveness of our algorithm.
{IEEEkeywords}
Distributed optimization, privacy preservation, differential privacy, sensitivity, gradient tracking
1 Introduction
[…]
With the widespread application of distributed optimization in various fields such as machine learning, power systems, and robotic networks, the growing sophistication of cyberattacks has brought security and privacy concerns to the forefront of research [1]. Compared to traditional centralized approaches, distributed optimization methods offer better privacy protection by avoiding the direct exchange of raw data, which may contain sensitive information. Nevertheless, studies have shown that the exchange of states and (estimated) gradients between agents in distributed optimization can still pose significant privacy risks. For instance, in distributed machine learning, attackers can accurately recover an agent’s raw data through its gradient [2]. In the distributed rendezvous problem, an agent’s state itself is its sensitive information [3]. Therefore, how to design distributed optimization methods that can protect the sensitive information of agents has become an urgent challenge.
In general, privacy-preserving distributed optimization algorithms can be categorized into three types: correlated randomness-based methods, encryption-based methods, and differential privacy-based methods. Correlated randomness-based methods achieve privacy protection by introducing temporally or spatially correlated randomness into optimization parameters [4, 5, 6, 7]. For example, [6] preserves the privacy of agents’ gradients by adding randomness to the stepsize and coupling weights during the initial iterations. By leveraging the robustness of dynamics, these methods can attain exact convergence. However, they necessitate an additional topological assumption to ensure privacy protection and cannot defend against external eavesdroppers. In contrast, encryption-based methods can effectively defend against external eavesdroppers by encrypting the transmitted information [8, 9, 10]. For example, in [10], agents use the Advanced Encryption Standard (AES) to encrypt transmitted information before sharing it with neighboring agents. However, encryption and decryption operations substantially increase the algorithm’s computational and communication overhead.
Due to the wide applicability, simplicity, and strong theoretical privacy guarantees, Differential Privacy (DP) has emerged as the de facto standard for privacy protection in recent years [11]. Several studies have integrated DP into distributed optimization by injecting random noise into the information transmitted between agents [12, 13, 14, 15, 16]. For example, [12] proposes a differentially private algorithm based on distributed projected gradient descent, under the assumption that the gradients of all individual cost functions are bounded. However, since distributed gradient descent relies on inexact gradients, its convergence rate remains sublinear, even without the addition of DP noise. In contrast, the gradient tracking algorithm can achieve a linear convergence rate with a constant stepsize. To this end, [13] presents a differentially private algorithm based on gradient tracking that achieves linear convergence. However, this approach assumes that different objective functions in two adjacent problems share the same Lipschitz constant and convexity parameter. On the other hand, both [12] and [13] reveal the inherent trade-off between privacy and accuracy in differentially private algorithms. To overcome this challenge, [14] proposes a distributed algorithm based on gradient tracking that ensures both DP and optimality by incorporating decaying weakening factors. However, to achieve a finite privacy budget as the number of iterations approaches infinity, this method requires the assumption that different objective functions in adjacent problems have consistent gradients within the neighborhood of the optimal solution of one problem. In addition, in the presence of persistent DP noise, gradient tracking techniques suffer from noise accumulation, which negatively impacts the algorithm’s optimization accuracy. To mitigate this issue, [15] proposes a differentially private distributed algorithm based on robust gradient tracking. However, its applicability is restricted to undirected graphs due to the use of doubly stochastic weight matrices. To address this limitation, [16] employs a row and a column stochastic weight matrices to develop a robust gradient tracking-based differentially private algorithm that extends applicability to directed graphs. [15] and [16] also reveal the inherent trade-off between privacy and accuracy in differentially private algorithms. Furthermore, [16] relaxes the assumption that the gradients of all individual cost functions are bounded, instead assuming that the distance between the gradients of different objective functions in adjacent problems is bounded.
Despite significant advancements in differentially private distributed optimization, several critical challenges remain. Firstly, existing differentially private algorithms lack design principles to enhance their accuracy, as well as theoretical validation for such improvements. Secondly, there is no direct analytical relationship between the noise parameter and the privacy budget. As a result, it becomes impractical to determine the required noise level for a given privacy budget before the algorithm’s execution, thus limiting real-world deployment. Thirdly, existing convergence analysis techniques become inapplicable in the presence of noise injection and diminishing stepsize. To address the limitations, this paper proposes corresponding solutions, which, along with the contributions of this work, are summarized as follows:
1) Enhancement of the accuracy of the differentially private distributed algorithm: We first reveal the relationship between the algorithm’s accuracy and sensitivity, demonstrating that lower sensitivity results in higher accuracy under the same privacy budget. This establishes a theoretical foundation for the design of the algorithm. Subsequently, we analyze the main factors affecting the sensitivity of the algorithm (Proposition 1). Building on this analysis, we propose an implicit gradient tracking-based differentially private distributed algorithm with lower sensitivity. Furthermore, we theoretically and empirically validate that the proposed algorithm achieves higher accuracy compared to existing methods [13, 15, 14, 16] under the same privacy budget.
2) Establishment of a closed-form expression for the noise parameter as a function of the privacy budget and reduction of the communication and computational burdens: For the proposed differentially private distributed algorithm, we derive a closed-form expression for the noise parameter as a function of the privacy budget. Based on this expression, it becomes feasible to directly compute the required noise parameter for any given privacy budget. Additionally, compared to [13, 15, 14, 16], the proposed algorithm reduces the number of communication variables by half, alleviating the communication burden. The number of injected noise is also halved, thereby reducing the computational burden of the algorithm.
3) Theoretical analysis of $\epsilon$-DP and convergence: Under a weak assumption on the gradients of different objective functions in adjacent problems, we prove that the proposed algorithm with appropriate parameters can achieve rigorous $\epsilon$-DP (Theorem 4.5). Additionally, under the assumption that the objective function is Lipschitz smooth and strongly convex, we establish the convergence of the proposed algorithm with diminishing stepsize and noise parameter, as well as an upper bound for its optimization accuracy (Theorem 5.17). The theoretical results reveal the trade-off between the algorithm’s privacy and accuracy. Based on these results, we further provide two tuning strategies for the algorithm’s parameters.
The remainder of this paper is organized as follows. Section 2 presents the problem formulation and provides relevant definitions of DP. Section 3 proposes a differentially private distributed algorithm with lower sensitivity. Sections 4 and 5 analyze the differential privacy guarantees and convergence of the proposed algorithm, respectively. Section 6 validates the effectiveness of the proposed algorithm through numerical simulations. Finally, Section 7 concludes the paper.
Notation: For any vector $\bm{v}\in\mathbb{R}^{p}$ and $1\leq a\leq\infty$, $\|\bm{v}\|_{a}$ represents the standard $L^{a}$-norm for $\bm{v}$. Without a subscript, $\|\bm{v}\|$ denotes $L^{2}$-norm of $\bm{v}$. For any matrix $M\in\mathbb{R}^{n\times p}$, $\|M\|_{a}$ represents the standard induced $L^{a}$-norm for $M$ such that $\|M\|_{a}\triangleq\sup_{\|x\|_{a}\not=0}\frac{\|Mx\|_{a}}{\|x\|_{a}}$. Without a subscript, $\|M\|$ denotes the Frobenius norm of $M$. $\bm{0},\bm{1}\in\mathbb{R}^{n}$ represent the vectors of all zeros and ones, respectively. $I\in\mathbb{R}^{n\times n}$ denotes the identity matrix. For a square matrix $M\in\mathbb{R}^{n\times n}$, $\rho(M)$ represents its spectral radius. Given a sequence of matrices $\{M_{k}\}$ with $M_{k}\in\mathbb{R}^{n\times n}$ and for $K_{1},K_{2}\geq 1$,
$\prod_{k=K_{1}}^{K_{2}}M_{k}=\left\{\begin{aligned} &M_{K_{2}}\cdots M_{K_{1}},&\text{if}\ K_{2}>K_{1},\\ &M_{K_{1}},&\text{if}\ K_{2}=K_{1},\\ &I,&\text{if}\ K_{2}<K_{1}.\end{aligned}\right.$
For the random variables $X$ and $Y$, $\mathbb{P}[X]$ represents the probability of $X$, $\mathbb{E}[X]$ denotes the expectation of $X$, and $\mathbb{E}[X|Y]$ is the conditional expectation of $X$ given $Y$. Given a constant $c>0$, $Lap(c)$ represents the Laplace distribution with the probability density function $p_{c}(x)\triangleq\frac{1}{2c}e^{-\frac{|x|}{c}}$. This distribution has a mean of zero and a variance of $2c^{2}$. For any $x,y\in\mathbb{R}$, it can be shown that $\frac{p_{c}(x)}{p_{c}(y)}\leq e^{\frac{|y-x|}{c}}$. If $x\sim Lap(c)$, $\mathbb{E}[x]=0$, $\mathbb{E}[|x|]=c$, and $\mathbb{E}[x^{2}]=2c^{2}$. For any matrix $M\in\mathbb{R}^{n\times p}$, $M\sim Lap(c)$ implies that each component $M_{ij}\sim Lap(c)$.
2 Problem Formulation
Consider a multi-agent system consisting of $n$ agents. Each agent $i$ possesses an individual cost function $f_{i}:\mathbb{R}^{p}\to\mathbb{R}$, which is only known by agent $i$ itself. All agents perform local computation and communication over an undirected network to collectively minimize:
$f(\bm{x})\triangleq\frac{1}{n}\sum_{i=1}^{n}f_{i}(\bm{x}),$ (1)
where $\bm{x}\in\mathbb{R}^{p}$ is the global decision variable and $f:\mathbb{R}^{p}\to\mathbb{R}$ is the average of all individual cost functions. Let $\bm{x}_{i}$ be a local copy of the decision variable for agent $i$. Denote the gradient of $f_{i}$ with respect to $\bm{x}_{i}$ as $\nabla f_{i}(\bm{x}_{i})$. For convenience, we further define $\mathbf{x}\triangleq\left[\bm{x}_{1},\cdots,\bm{x}_{n}\right]^{T}\in\mathbb{R}^{n\times p}$ and $\nabla F(\mathbf{x})\triangleq\left[\nabla f_{1}(\bm{x}_{1}),\cdots,\nabla f_{n}(\bm{x}_{n})\right]^{T}\in\mathbb{R}^{n\times p}$.
We model an undirected communication network using the graph $\mathcal{G}\triangleq(\mathcal{N},\mathcal{E})$, where $\mathcal{N}\triangleq\{1,2,\cdots,n\}$ represents the set of agents and $\mathcal{E}\subseteq\mathcal{N}\times\mathcal{N}$ denotes the set of edges. $(i,j)\in\mathcal{E}$ if and only if agents $i$ and $j$ can send information to each other. The set of agents that can communicate with agent $i$ is defined as $\mathcal{N}_{i}\triangleq\left\{j\in\mathcal{N}:(i,j)\in\mathcal{E}\right\}$. Moreover, we employ a nonnegative and symmetric matrix $W\in\mathbb{R}^{n\times n}$ to represent the inter-agent weights. For any $i,j\in\mathcal{N}$, $W_{ij}>0$ if $i=j$ or $(i,j)\in\mathcal{E}$. For all other $(i,j)$, $W_{ij}=0$. We define $\mathcal{G}_{W}\triangleq(\mathcal{N},\mathcal{E}_{W})$ as the graph induced by the weight matrix $W$, where $(i,j)\in\mathcal{E}_{W}$ if and only if $i\not=j$ and $W_{ij}>0$.
Throughout the paper, we make the following assumptions:
Assumption 1
For each $i\in\mathcal{N}$, the individual cost function $f_{i}$ is $L$-smooth and $\mu$-strongly convex, i.e., for any $\bm{x},\bm{x}^{\prime}\in\mathbb{R}^{p}$,
$\|\nabla f_{i}(\bm{x})-\nabla f_{i}(\bm{x}^{\prime})\|\leq L\|\bm{x}-\bm{x}^{\prime}\|,$
$f_{i}(\bm{x}^{\prime})\geq f_{i}(\bm{x})+\nabla f_{i}(\bm{x})^{T}(\bm{x}^{\prime}-\bm{x})+\frac{\mu}{2}\|\bm{x}^{\prime}-\bm{x}\|^{2}.$
Assumption 2
The nonnegative weight matrix $W$ is symmetric and doubly-stochastic, satisfying $W=W^{T}$, $\bm{1}^{T}W=\bm{1}^{T}$, and $W\bm{1}=\bm{1}$. Its induced graph $\mathcal{G}_{W}$ is connected, that is, there exists a path between any two agents.
Next, we introduce some notations. We describe the distributed optimization problem $\mathcal{P}$ in (1) using the tuple $(\mathcal{X},\mathcal{F},f,\mathcal{G}_{W})$, where $\mathcal{X}=\mathbb{R}^{p}$ is the domain of optimization. $\mathcal{F}\subseteq\{\mathcal{X}\to\mathbb{R}\}$ is a set of real-valued cost functions. $f(\bm{x})\triangleq\frac{1}{n}\sum_{i=1}^{n}f_{i}(\bm{x})$ with $f_{i}\in\mathcal{F}$ for each $i\in\mathcal{N}$. $\mathcal{G}_{W}$ is a communication graph induced by the weight matrix $W$. We represent a distributed optimization algorithm for solving the problem $\mathcal{P}$ by a mapping $Alg\left(\mathcal{P},\mathbf{x}(0)\right):\{\mathcal{P},\mathbf{x}(0)\}\to\mathcal{Z}$, where $\mathbf{x}(0)\in\mathcal{X}^{n}$ is the initial state of all agents. $\mathcal{Z}=\{\mathbf{z}(0),\mathbf{z}(1),\cdots\}\in\mathbb{O}$ is the sequence of shared information of all agents, also known as the observation sequence, and $\mathbb{O}$ denotes the set of all possible observation sequences. Furthermore, under a distributed optimization algorithm $Alg$, for $\mathcal{P}$, $\mathbf{x}(0)$, and $\mathcal{Z}$, we denote the algorithm’s internal states as $Sta(\mathcal{P},\mathbf{x}(0),\mathcal{Z})$.
With the above notations, we are ready to present the definitions of adjacency for two distributed optimization problems, and of $\epsilon$-DP and sensitivity for a distributed algorithm.
Definition 1
(Adjacency) Two distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{\prime}$ are adjacent, if the following hold:
  1. 1.
    $\mathcal{X}=\mathcal{X}^{\prime}$, $\mathcal{F}=\mathcal{F}^{\prime}$, and $\mathcal{G}_{W}=\mathcal{G}_{W}^{\prime}$.
  2. 2.
    There exists an $i_{0}\in\mathcal{N}$ such that $f_{i_{0}}\not=f_{i_{0}}^{\prime}$ and for all $j\not=i_{0}$, $f_{j}=f_{j}^{\prime}$.
  3. 3.
    The distance between the gradients of $f_{i_{0}}$ and $f_{i_{0}}^{\prime}$ is bounded by $\delta$ on $\mathcal{X}$, i.e., for any $\bm{x}\in\mathcal{X}$, $\|\nabla f_{i_{0}}(\bm{x})-\nabla f_{i_{0}}^{\prime}(\bm{x})\|_{1}\leq\delta$.
From Definition 1, two distributed optimization problems are adjacent if only one agent modifies its cost function while all other parameters remain unchanged. Furthermore, we require that the distance between the gradients of two different cost functions is bounded. To the best of our knowledge, this condition is the most relaxed in the existing literature, and a detailed discussion is provided in the following remark.
Remark 1
Regarding the condition on different cost functions in adjacent problems, [12] and [15] require that the gradients of all cost functions be bounded, i.e., for any $\bm{x}\in\mathcal{X}$ and $f_{i}\in\mathcal{F}$, $\|\nabla f_{i}(\bm{x})\|\leq\delta_{1}$ with $\delta_{1}>0$. In contrast, our requirement in Definition 1 is more relaxed, allowing for a broader class of cost functions. For example, the functions $f_{i_{0}}(\bm{x})=\frac{1}{2}\bm{x}^{T}\bm{x}$ and $f_{i_{0}}^{\prime}(\bm{x})=\frac{1}{2}\bm{x}^{T}\bm{x}+\frac{\delta}{p}\bm{x}$ satisfy our requirement, but their gradients are unbounded on $\mathbb{R}^{p}$. Furthermore, [13] assumes that $f_{i_{0}}$ and $f_{i_{0}}^{\prime}$ share the same Lipschitz and convexity constants. [14] assumes that $f_{i_{0}}$ and $f_{i_{0}}^{\prime}$ have identical gradients in the neighborhood of the solution of $\mathcal{P}$. These assumptions impose significantly stricter constraints compared to ours.
Definition 2
($\epsilon$-DP) Given an $\epsilon>0$, a distributed optimization algorithm $Alg$ is $\epsilon$-differentially private if for any two adjacent distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{\prime}$, any initial state $\mathbf{x}(0)\in\mathcal{X}^{n}$, and any set of the observation sequences $\mathcal{O}\subseteq\mathbb{O}$,
$\mathbb{P}\left[Alg\left(\mathcal{P},\mathbf{x}(0)\right)\in\mathcal{O}\right]\leq e^{\epsilon}\mathbb{P}\left[Alg\left(\mathcal{P}^{\prime},\mathbf{x}(0)\right)\in\mathcal{O}\right],$ (2)
where the probability $\mathbb{P}$ is taken over the randomness of the algorithm.
By Definition 2, a distributed algorithm with $\epsilon$-DP ensures that the distribution of any agent’s shared information remains nearly indistinguishable between adjacent problems. In other words, an adversary who has obtained the observation sequence cannot identify the individual cost function of any one agent with significant probability. In addition, a smaller $\epsilon$ indicates a higher level of privacy.
Definition 3
(Sensitivity) At each iteration $k$, for any two adjacent distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{\prime}$ and any initial state $\mathbf{x}(0)\in\mathcal{X}^{n}$, the sensitivity of the algorithm is
$\Delta(k)\triangleq\sup_{\mathcal{Z}\in\mathbb{O}}\sup_{\begin{subarray}{c}\mathbf{x}(k)\in Sta(\mathcal{P},\mathbf{x}(0),\mathcal{Z})\\ \mathbf{x}^{\prime}(k)\in Sta(\mathcal{P}^{\prime},\mathbf{x}(0),\mathcal{Z})\end{subarray}}\|\mathbf{x}(k)-\mathbf{x}^{\prime}(k)\|_{1}.$ (3)
The sensitivity $\Delta(k)$ characterizes the maximum deviation in the algorithm’s states at iteration $k$ under adjacent problems. More importantly, the algorithm’s sensitivity directly influences the noise variance required to achieve DP. According to Lemma 2 in [12] (or see Lemma 4.4 in Section 4), the sensitivity of the algorithm is positively correlated with the noise variance under a fixed privacy budget, i.e., lower sensitivity necessitates a smaller noise variance. Conversely, the algorithm’s optimization accuracy is inversely proportional to the noise variance, i.e., the smaller the noise variance, the higher the algorithm’s accuracy [16]. In summary, reducing the sensitivity of the algorithm can enhance its accuracy. This insight provides a clear guideline for designing differentially private algorithms with enhanced accuracy: minimizing the algorithm’s sensitivity as much as possible.
3 Algorithm Design
In this section, we propose a differentially private distributed algorithm with lower sensitivity. Additionally, we conduct a comparative analysis of our algorithm against related methods.
3.1 Algorithm Design
As discussed in the previous section, the key to designing a differentially private algorithm with enhanced accuracy lies in reducing the algorithm’s sensitivity. Thus, identifying the main factors that affect the algorithm’s sensitivity becomes crucial. Generally, distributed optimization algorithms consist of two main steps: a consensus step and an optimization step. In the consensus step, each agent combines its state with those of neighboring agents to minimize the consensus error related to other agents’ states. In the optimization step, each agent updates its state along the gradient direction to reduce the optimization error associated with Problem (1). To achieve DP, each agent needs to inject random noise into its shared state, which we refer to as the noisy state. However, when designing a differentially private distributed algorithm, two issues arise:
  1. 1.
    In the consensus step, should agents use their true states or noisy states?
  2. 2.
    In the optimization step, should agents compute gradients at true states or noisy states?
We argue that both the design choices mentioned above and the amount of shared information influence the algorithm’s sensitivity. The key findings are summarized as follows:
Proposition 1
Consider two differentially private distributed algorithms (Alg 1 and Alg 2). Let $\Delta_{1}(k)$ and $\Delta_{2}(k)$ be the sensitivities of the two algorithms, respectively. Under Definition 1, the following claims hold:
  1. (i)
    Alg 1: agents use their noisy states in the consensus step. Alg 2: agents use their true states in the consensus step. Then, $\Delta_{1}(k)\leq\Delta_{2}(k)$ for all $k\geq 1$.
  2. (ii)
    Alg 1: agents compute gradients based on noisy states. Alg 2: agents compute gradients based on true states. Then, $\Delta_{1}(k)\leq\Delta_{2}(k)$ for all $k\geq 1$.
  3. (iii)
    Alg 1: agents exchange one variable $\bm{x}$. Alg 2: agents exchange two variables $\bm{x}$ and $\bm{y}$. Then, $\Delta_{1}(k)\leq\Delta_{2}(k)$ for all $k\geq 1$.
Proof 3.1.
The proof is provided in Appendix 8.
Algorithm 1 Differentially Private Distributed Algorithm with Lower Sensitivity
0:  Initialization $\bm{x}_{i}(0)\in\mathbb{R}^{p}$ and $\bm{y}_{i}(0)=\bm{0}$, the weight matrix $W$, the stepsize sequence $\{\alpha_{k}\}_{k=1}^{\infty}$, $\beta>0$.
0:
1:   Each agent $i$ injects noise $\bm{\xi}_{i}(k)$ into $\bm{x}_{i}(k-1)$.
2:   Each agent $i$ sends $\bm{z}_{i}(k)=\bm{x}_{i}(k-1)+\bm{\xi}_{i}(k)$ to $j\in\mathcal{N}_{i}$ and receives $\bm{z}_{j}(k)$ from $j\in\mathcal{N}_{i}$.
3:   Each agent $i$ executes the following updates:
$\bar{\bm{z}}_{i}(k)=\sum_{j=1}^{n}W_{ij}\bm{z}_{j}(k),$ (4a)
$\bm{y}_{i}(k)=\bm{y}_{i}(k-1)+\beta\left(\bm{z}_{i}(k)-\bar{\bm{z}}_{i}(k)\right),$ (4b)
$\bm{x}_{i}(k)=\bar{\bm{z}}_{i}(k)-\alpha_{k}\left(\bm{y}_{i}(k)+\nabla f_{i}\left(\bm{z}_{i}(k)\right)\right).$ (4c)
Building on the design principles in Proposition 1, we propose a differentially private distributed algorithm with lower sensitivity, as detailed in Algorithm 1. In this algorithm, $\bm{y}_{i}$ serves as an auxiliary variable that compensates for the error between the local gradient and the average gradient, thereby accelerating convergence. Its principle will be discussed in the next subsection. To achieve $\epsilon$-DP, Algorithm 1 injects additive noise $\bm{\xi}_{i}$ into the shared information of agent $i$ at every iteration $k$. Notably, in Algorithm 1, each agent $i$ utilizes its noisy state $\bm{z}_{i}$ in the consensus step (4a), computes the gradient $\nabla f_{i}\left(\cdot\right)$ at its noisy state $\bm{z}_{i}$ in the optimization step (4c), and shares only the variable $\bm{z}_{i}$ to its neighboring agents, without disclosing the auxiliary variable $\bm{y}_{i}$. These designs adhere to Proposition 1 and ensure that Algorithm 1 achieves lower sensitivity.
Let $\bm{\xi}(k)\triangleq[\bm{\xi}_{1}(k),\bm{\xi}_{2}(k),\cdots,\bm{\xi}_{n}(k)]^{T}\in\mathbb{R}^{n\times p}$. The injected random noise satisfies the following assumption:
Assumption 3
The random noise $\bm{\xi}(k)$ is independent across dimensions $p$, agents $i$, and iterations $k$. Moreover, it follows a zero-mean Laplace distribution with parameter $\nu_{k}$, i.e., $\bm{\xi}(k)\sim Lap(\nu_{k})$.
For the stepsize $\alpha_{k}$ and the noise parameter $\nu_{k}$, we choose them as exponentially decaying sequences:
$\alpha_{k}=\gamma q_{1}^{k-1},\quad\nu_{k}=\frac{\gamma\delta q_{2}}{\epsilon(q_{2}-q_{1})}q_{2}^{k-1},$ (5)
where $\gamma$ is the initial stepsize parameter with $\gamma\beta\leq 1$, $q_{1}\in(0,1)$ is the stepsize decay parameter, $\delta$ is the bound on the gradient distance in Definition 1, $q_{2}\in(q_{1},1)$ is the noise decay parameter, and $\epsilon>0$ is the privacy budget.
Table 1: Comparison of Algorithm 1 with Existing Algorithms.
Algorithm Closed-form expression Number of Communica-
for noise w.r.t. budget $\epsilon$ added noise tion
[13] $2p$ $2p$
[15] $2p$ $2p$
[14] $2p$ $2p$
[16] $2p$ $2p$
Our method $p$ $p$
Algorithm 1 involves four tunable parameters $\gamma$, $\beta$, $q_{1}$, and $q_{2}$. For these parameters, we provide two tuning strategies in Section 5. Using parameters (5), Algorithm 1 achieves $\epsilon$-DP, convergence, and a guaranteed level of accuracy. These properties will be analyzed in detail in Sections 4 and 5. Moreover, we summarize the advantages of Algorithm 1 over existing state-of-the-art algorithms [13, 15, 14, 16] in Table 1, with detailed discussions provided in the following remarks.
Remark 3.2.
For Algorithm 1, given any privacy budget $\epsilon>0$, the noise parameter $\nu_{k}$ can be easily obtained using Eq. (5). This is more convenient and practical compared to algorithms in [13, 15, 14, 16], in which there are no closed-form expression for the noise parameter with respect to the privacy budget.
Remark 3.3.
In the proposed algorithm, only variable $\bm{x}$ is shared among agents, while the auxiliary variable $\bm{y}$ remains local and is not exchanged. Compared to existing algorithms [13, 15, 14, 16], the proposed algorithm reduces both the number of communication variables and the required injected noise by half. As a result, it significantly reduces communication and computational burdens.
3.2 Related Methods
To get a better insight into the update rule of our algorithm, we compare it with the Distributed Gradient Descent (DGD) and Gradient Tracking (GT) methods. Notably, [12] and [14] develop differentially private distributed algorithms based on the (projected) DGD method. The update rule for DGD is
$\mathbf{x}(k+1)=W\mathbf{x}(k)-\alpha\nabla F(\mathbf{x}(k)).$ (6)
The DGD method with a constant stepsize cannot converge to the optimal solution. To illustrate it, assume that $\mathbf{x}(k)$ converges to a consistent limit $\mathbf{x}(\infty)$. Since $W\bm{1}=\bm{1}$, it follows that $\mathbf{x}(\infty)=W\mathbf{x}(\infty)$. Taking the limit as $k\to\infty$ on both sides of (6), we obtain $\nabla F(\mathbf{x}(\infty))=\bm{0}$. This means that all local cost functions $f_{i}$ attain their minimum at the same $\bm{x}_{i}(\infty)$, which is generally impossible. Thus, the DGD method’s stepsize must be decaying to ensure exact convergence. However, the decaying stepsize results in slow convergence both in theory and practice [17].
The GT method overcomes the above limitation by introducing an auxiliary variable $\mathbf{y}$, which enables exact convergence with a constant stepsize. Note that the differentially private distributed algorithms in [13, 15, 14, 16] are based on the GT method or its variants. The update rule for GT is given by:
$\mathbf{x}(k+1)$ $=W\mathbf{x}(k)-\alpha\mathbf{y}(k),$ (7a)
$\mathbf{y}(k+1)$ $=W\mathbf{y}(k)+\nabla F(\mathbf{x}(k+1))-\nabla F(\mathbf{x}(k)),$ (7b)
where $\mathbf{y}(0)=\nabla F(\mathbf{x}(0))$. To demonstrate how GT achieves exact convergence, we first substitute (7a) into (7b) to derive the formula of $\mathbf{x}(k+2)-\mathbf{x}(k+1)$. Next, combining the first update $\mathbf{x}(1)=W\mathbf{x}(0)-\alpha\nabla F(\mathbf{x}(0))$ with $\mathbf{x}(2)-\mathbf{x}(1),\cdots,\mathbf{x}(k+1)-\mathbf{x}(k)$ yields:
$\mathbf{x}(k+1)=$ $\ W\mathbf{x}(k)-\alpha\nabla F(\mathbf{x}(k))+\sum_{l=1}^{k}\left(W-I\right)\mathbf{x}(l)$ (8)
$+\sum_{l=0}^{k-1}(W-W^{2})\mathbf{x}(l).$
In comparison with the DGD method, the GT method introduces two additional terms. Taking the limit as $k\to\infty$ on both sides of (8), we obtain $\alpha\nabla F(\mathbf{x}(\infty))=\sum_{l=1}^{\infty}\left(W-I\right)\mathbf{x}(l)+\sum_{l=0}^{\infty}(W-W^{2})\mathbf{x}(l)$. Since $\bm{1}^{T}(W-I)=0$ and $\bm{1}^{T}(W-W^{2})=0$, we obtain $\bm{1}^{T}\nabla F(\mathbf{x}(\infty))=0$, which corresponds to the objective of finding a point that minimizes the average of $\sum_{i=1}^{n}f_{i}$.
For Algorithm 1, to better illustrate its fundamental mechanism, we omit the noise term and replace the decaying stepsize $\alpha_{k}$ with a fixed stepsize $\alpha$, yielding the following update rule:
$\mathbf{y}(k+1)$ $=\mathbf{y}(k)+\beta(I-W)\mathbf{x}(k),$ (9a)
$\mathbf{x}(k+1)$ $=W\mathbf{x}(k)-\alpha\left(\mathbf{y}(k+1)+\nabla F(\mathbf{x}(k))\right),$ (9b)
where $\mathbf{y}(0)=\bm{0}$. From (9a), we can deduce that $\mathbf{y}(k+1)=\sum_{l=0}^{k}\beta(I-W)\mathbf{x}(l)$. Substituting this into (9b) gives:
$\mathbf{x}(k+1)=W\mathbf{x}(k)-\alpha\nabla F(\mathbf{x}(k))+\alpha\beta\sum_{l=0}^{k}\left(W-I\right)\mathbf{x}(l).$ (10)
In comparison with the GT method (8), when $\alpha\beta=1$, our algorithm introduces the term $(W-I)\mathbf{x}(0)$ but eliminates the term $\sum_{l=0}^{k-1}(W-W^{2})\mathbf{x}(l)$. Taking the limit as $k\to\infty$ on both sides of (10), we obtain $\nabla F(\mathbf{x}(\infty))=\beta\sum_{l=0}^{\infty}(W-I)\mathbf{x}(l)$. Since $\mathbf{1}^{T}(W-I)=0$, it follows that $\mathbf{1}^{T}\nabla F(\mathbf{x}(\infty))=0$. Roughly speaking, our algorithm, similar to the GT method, can achieve exact convergence with a constant stepsize. More importantly, in our algorithm, neighboring agents only share the variable $\bm{x}$, whereas in the GT method, both $\bm{x}$ and the auxiliary variable $\bm{y}$ need to be exchanged. Thus, our algorithm reduces the communication burden by half compared to the GT method. On the other hand, in the GT method, the auxiliary variable $\mathbf{y}$, satisfying $\frac{1}{n}\mathbf{1}^{T}\mathbf{y}(k)=\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{x}(k))$ for any $k\geq 0$, is used to track the average gradient. In our algorithm, the auxiliary variable $\mathbf{y}$ satisfies $\frac{1}{n}\mathbf{1}^{T}(\mathbf{y}(k)+\nabla F(\mathbf{x}(k)))=\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{x}(k))$ for any $k\geq 0$, which can be interpreted as tracking the difference between the average gradient and the local gradient.
4 Differential Privacy
Table 2: Comparison of Sensitivity of Algorithm 1 with Existing Algorithms.
Algorithm Sensitivity
[13] Not lower than $\bm{(2c_{1}+2)(\delta+c_{2})}$ with $c_{1},c_{2}>0$.
[15] Not lower than $\bm{(2c_{3}+1)(\delta+c_{4})\alpha_{1}(k)}$ with $c_{3},c_{4}>0$ and $\sum_{k=1}^{\infty}\alpha_{1}(k)<\infty$.
[14] Not lower than $\bm{(2-\alpha_{2}(k))(\delta+c_{5})(\alpha_{3}(k)+1)}$ with $\sum_{k=1}^{\infty}\alpha_{2}(k)=\infty$, $\lim_{k\to\infty}\alpha_{2}(k)=0$, $\sum_{k=1}^{\infty}\alpha_{3}(k)=\infty$, $\lim_{k\to\infty}\alpha_{3}(k)=0$, and $c_{5}>0$.
[16] Not lower than $\bm{3(\delta+c_{6})\alpha_{4}(k)}$ with $c_{6}>0$ and $\sum_{k=1}^{\infty}\alpha_{4}(k)<\infty$.
Our method No more than $\bm{\delta\alpha_{k}}$ with $\sum_{k=1}^{\infty}\alpha_{k}<\infty$.
In this section, we show that Algorithm 1 with the parameter settings (5) guarantees $\epsilon$-DP. We then validate the advantages of Algorithm 1 in terms of accuracy.
The following lemma provides a sufficient condition for Algorithm 1 to guarantee $\epsilon$-DP.
Lemma 4.4.
([12, Lemma 2]) If each agent $i$ adds a noise vector $\bm{\xi}_{i}(k)$ under Assumption 3 such that $\sum_{k=1}^{\infty}\frac{\Delta(k)}{\nu_{k}}\leq\epsilon$, then Algorithm 1 is $\epsilon$-differentially private.
In the following theorem, we first give an upper bound on the sensitivity of Algorithm 1. Then, based on Lemma 4.4, we prove that Algorithm 1, with the parameter settings given in (5), is $\epsilon$-differentially private.
Theorem 4.5.
Under Definitions 1 and 3, the sensitivity of Algorithm 1 satisfies $\Delta(k)\leq\delta\alpha_{k}$. Given any $\epsilon>0$, if the random noise $\bm{\xi}(k)\sim Lap(\nu_{k})$ satisfies Assumption 3, and the stepsize $\alpha_{k}$ and the noise parameter $\nu_{k}$ are chosen according to (5), then Algorithm 1 guarantees $\epsilon$-DP.
Proof 4.6.
With the same initial state $\mathbf{x}(0)$ and observation sequence $\mathcal{Z}$, for any two adjacent problems $\mathcal{P}$ and $\mathcal{P}^{\prime}$, we have $\bm{z}_{i}(k)=\bm{z}_{i}^{\prime}(k)$ for all $i\in\mathcal{N}$ and all $k\geq 1$. Based on (4a), (4b), and $\bm{y}_{i}(0)=\bm{y}_{i}^{\prime}(0)$, it follows that $\bm{y}_{i}(k)=\bm{y}_{i}^{\prime}(k)$ for all $i\in\mathcal{N}$ and all $k\geq 1$. From Definition 1, we know that $f_{j}\not=f_{j}^{\prime}$ if and only if $j=i_{0}$. In light of (4c), it follows that $\bm{x}_{j}(k)=\bm{x}_{j}^{\prime}(k)$ for all $j\in\mathcal{N}$, $j\not=i_{0}$ and all $k\geq 1$. Therefore, the sensitivity of Algorithm 1 at iteration $k$ is bounded as follows:
$\|\mathbf{x}(k)-\mathbf{x}^{\prime}(k)\|_{1}$ $=\|\bm{x}_{i_{0}}(k)-\bm{x}_{i_{0}}^{\prime}(k)\|_{1}$ (11)
$=\alpha_{k}\|\nabla f_{i_{0}}(\bm{z}_{i_{0}}(k))-\nabla f_{i_{0}}^{\prime}(\bm{z}_{i_{0}}^{\prime}(k))\|_{1}$
$=\alpha_{k}\|\nabla f_{i_{0}}(\bm{z}_{i_{0}}(k))-\nabla f_{i_{0}}^{\prime}(\bm{z}_{i_{0}}(k))\|_{1}$
$\leq\delta\alpha_{k}.$
It follows from Lemma 4.4 that Algorithm 1 is $\epsilon$-differentially private if $\sum_{k=1}^{\infty}\frac{\delta\alpha_{k}}{\nu_{k}}\leq\epsilon$. When $\alpha_{k}=\gamma q_{1}^{k-1}$ and $\nu_{k}=\frac{\gamma\delta q_{2}}{\epsilon(q_{2}-q_{1})}q_{2}^{k-1}$, and given that $\gamma>0$, $q_{1}\in(0,1)$, $q_{2}\in(q_{1},1)$, we have
$\sum_{k=1}^{\infty}\frac{\delta\alpha_{k}}{\nu_{k}}=\frac{\epsilon(q_{2}-q_{1})}{q_{2}}\sum_{k=1}^{\infty}\left(\frac{q_{1}}{q_{2}}\right)^{k-1}=\epsilon.$ (12)
Therefore, Algorithm 1 with the parameter settings given in (5) guarantees $\epsilon$-DP.
Remark 4.7.
In Table 2, we compare the sensitivity of differentially private distributed algorithms under the same definitions and assumptions outlined in this paper. As shown, our algorithm demonstrates lower sensitivity compared to algorithms in [13, 15, 14, 16]. This advantage stems from its adherence to the design principles outlined in Proposition 1, whereas other algorithms did not consider the sensitivity. According to Lemma 4.4, for a fixed privacy budget, the algorithm’s sensitivity is positively correlated with the noise variance. That is, lower sensitivity requires smaller noise variance. This leads to possible improvement of the optimization accuracy. Therefore, under the same privacy budget, our algorithm can achieve higher optimization accuracy than existing state-of-the-art algorithms.
5 Convergence
In this section, we first present some fundamental lemmas, followed by the convergence proof of Algorithm 1 and analysis of the optimization accuracy. We show clearly the trade-off between privacy and accuracy and provide guidance for tuning the algorithm’s parameters.
5.1 Preliminary Analysis and Fundamental Lemmas
We rewrite Algorithm 1 in an augmented form as follows:
$\mathbf{y}(k)$ $=\mathbf{y}(k-1)+\beta\left(I-W\right)\mathbf{z}(k),$ (13a)
$\mathbf{x}(k)$ $=W\mathbf{z}(k)-\alpha_{k}\left(\mathbf{y}(k)+\nabla F(\mathbf{z}(k))\right),$ (13b)
where $\mathbf{y}(0)=\mathbf{0}$ and $\mathbf{z}(k)=\mathbf{x}(k-1)+\bm{\xi}(k)$. Given that $W$ is doubly-stochastic, it follows from (13a) that $\frac{1}{n}\mathbf{1}^{T}\mathbf{y}(k)=0$ for all $k\geq 1$. Define $\bar{\bm{x}}(k)\triangleq\frac{1}{n}\mathbf{1}^{T}\mathbf{x}(k)$. From (13b), we obtain
$\bar{\bm{x}}(k)$ $=\bar{\bm{x}}(k-1)-\frac{\alpha_{k}}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))+\frac{1}{n}\mathbf{1}^{T}\bm{\xi}(k).$ (14)
According to Assumption 1, Problem (1) has a unique optimal solution, denoted as $\bm{x}^{*}$. We further have
$\bar{\bm{x}}(k+1)-\bm{x}^{*}=\bar{\bm{x}}(k)-\frac{\alpha_{k+1}}{n}\mathbf{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))-\bm{x}^{*}$ (15)
$\ -\frac{\alpha_{k+1}}{n}\mathbf{1}^{T}\left(\nabla F(\mathbf{z}(k+1))-\nabla F(\mathbf{1}\bar{\bm{x}}(k))\right)+\frac{1}{n}\mathbf{1}^{T}\bm{\xi}(k+1).$
Using (13a) to represent $\mathbf{y}(k+1)$ in terms of $\mathbf{y}(k)$ and $\mathbf{z}(k+1)$, and next eliminating $\mathbf{y}(k)$ via (13b), we get
$\mathbf{y}$ $(k+1)=\ \frac{1}{\alpha_{k}}W\mathbf{x}(k-1)+\Big{(}(\beta-\frac{1}{\alpha_{k}})I-\beta W\Big{)}\mathbf{x}(k)$ (16)
$-\nabla F(\mathbf{z}(k))+\beta(I-W)\bm{\xi}(k+1)+\frac{1}{\alpha_{k}}W\bm{\xi}(k).$
Then, it follows from (13b) and (14) that
$\quad\mathbf{x}(k+1)-\mathbf{1}\bar{\bm{x}}(k+1)$ (17)
$=W\mathbf{x}(k)-\alpha_{k+1}\mathbf{y}(k+1)-\alpha_{k+1}\nabla F(\mathbf{z}(k+1))-\mathbf{1}\bar{\bm{x}}(k)$
$\quad+\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k+1))+W\bm{\xi}(k+1)-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k+1)$
$=\tilde{W}(k)\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)+\frac{\alpha_{k+1}}{\alpha_{k}}W(\mathbf{x}(k)-\mathbf{x}(k-1))$
$\quad-\alpha_{k+1}(\nabla F(\mathbf{z}(k+1))-\nabla F(\mathbf{z}(k)))$
$\quad+\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k+1))-\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)$
$\quad+\Big{(}(1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\Big{)}\bm{\xi}(k+1),$
where $\tilde{W}(k)\triangleq(\frac{\alpha_{k+1}}{\alpha_{k}}-\alpha_{k+1}\beta)I+(1-\frac{\alpha_{k+1}}{\alpha_{k}}+\alpha_{k+1}\beta)W$ and the second equality follows from (16).
Remark 5.8.
Given that $\alpha_{k}=\gamma q_{1}^{k-1}$ with $\gamma>0$, $q_{1}\in(0,1)$, and $\gamma\beta\leq 1$, we have $\frac{\alpha_{k+1}}{\alpha_{k}}-\alpha_{k+1}\beta=q_{1}-\gamma\beta q_{1}^{k}\in[0,1)$ for all $k\geq 1$. Since $W$ is doubly-stochastic, we conclude that $\tilde{W}(k)$ is also a doubly-stochastic matrix for all $k\geq 1$. Let $\sigma_{k}\triangleq\|\tilde{W}(k)-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|$. Since the graph induced by $\tilde{W}(k)$ is connected, we have $\sigma_{k}\in(0,1)$ for all $k\geq 1$.
Using (13b) and (16) again, we have
$\quad\mathbf{x}(k+1)-\mathbf{x}(k)$ (18)
$=W\mathbf{z}(k+1)-\alpha_{k+1}\mathbf{y}(k+1)-\alpha_{k+1}\nabla F(\mathbf{z}(k+1))-\mathbf{x}(k)$
$=\frac{\alpha_{k+1}}{\alpha_{k}}W(\mathbf{x}(k)-\mathbf{x}(k-1))$
$\quad+\Big{(}1-\frac{\alpha_{k+1}}{\alpha_{k}}+\alpha_{k+1}\beta\Big{)}(W-I)\mathbf{x}(k)$
$\quad-\alpha_{k+1}(\nabla F(\mathbf{z}(k+1))-\nabla F(\mathbf{z}(k)))$
$\quad+\left((1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I\right)\bm{\xi}(k+1)-\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k).$
Let $\mathcal{M}_{k}$ be the $\sigma$-algebra generated by $\{\bm{\xi}(1),\cdots,\bm{\xi}(k-1)\}$. Define $\mathbb{E}\left[\cdot|\mathcal{M}_{k}\right]$ as the conditional expectation given $\mathcal{M}_{k}$. In the following, we present some auxiliary lemmas that will be used in the analysis.
Lemma 5.9.
([18, Lemma 2]) Under Assumption 1, the following inequalities hold:
$\|\mathbf{1}^{T}\nabla F(\mathbf{x}(k))-\mathbf{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))\|\leq\sqrt{n}L\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|,$
$\|\mathbf{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))\|\leq nL\|\bar{\bm{x}}(k)-\bm{x}^{*}\|.$
In addition, when $\alpha^{\prime}\leq\frac{2}{\mu+L}$, it follows that
$\|\bar{\bm{x}}(k)-\frac{\alpha^{\prime}}{n}\mathbf{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))-\bm{x}^{*}\|\leq(1-\alpha^{\prime}\mu)\|\bar{\bm{x}}(k)-\bm{x}^{*}\|.$
Lemma 5.10.
The following inequalities hold for all $k\geq 1$:
  1. (a)
    $\mathbb{E}\left[\langle\mathbf{z}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]\leq 2np\nu_{k}^{2}$.
  2. (b)
    $\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]\leq 2npL\nu_{k}^{2}$.
  3. (c)
    $\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k+1)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]\leq npL(2+np+2L\alpha_{k})\nu_{k}^{2}$.
  4. (d)
    $\mathbb{E}\left[\langle\mathbf{x}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]\leq 2np\left(3+L\alpha_{k}\right)\nu_{k}^{2}$.
  5. (e)
    $\mathbb{E}\left[\langle\mathbf{1}\bar{\bm{x}}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]\leq 2np(1+L\alpha_{k})\nu_{k}^{2}$.
  6. (f)
    $\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]\leq 2np\nu_{k+1}^{2}+\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$.
  7. (g)
    $\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{z}(k)\|^{2}|\mathcal{M}_{k}\right]\leq 2np\nu_{k+1}^{2}+2np(7+2L\alpha_{k})\nu_{k}^{2}+\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$.
Proof 5.11.
The proof can be found in Appendix 9.1.
Lemma 5.12.
Let $M\in\mathbb{R}^{n\times n}$ be a nonnegative, irreducible matrix with $n=2$ or $n=3$, and let $M_{ii}<\lambda^{*}$ for some $\lambda^{*}>0$. Then, $\rho(M)<\lambda^{*}$ if and only if $\det(\lambda^{*}I-M)>0$.
Proof 5.13.
The proof for the case $n=3$ can be found in Lemma 5 in [19]. The proof for the case $n=2$ is similar to that of $n=3$.
5.2 Main Results
Before presenting the main results, we first state some useful equations and inequalities. From the parameter settings in (5), we obtain $\frac{\alpha_{k+1}}{\alpha_{k}}=q_{1}<1$ and $\nu_{k+1}\leq\nu_{k}$. According to Assumption 2 and Remark 5.8, we know that both $W$ and $\tilde{W}(k)$ are symmetric doubly-stochastic matrices. This implies that the eigenvalues of $W$ and $\tilde{W}(k)$ lie within $[-1,1]$. Since the induced graphs are connected, both $W$ and $\tilde{W}(k)$ have a unique largest eigenvalue equal to 1, with eigenvector $\bm{1}$. Thus, we have $\|W\|=1$, $\|\tilde{W}(k)\|=1$, and $\|\tilde{W}(k)-I\|\leq 2$. Since $10<\alpha_{k+1}\beta=\gamma\beta q_{1}^{k}<$ 1, we have $\|(1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I\|\leq 3$ and $\|(1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|\leq 3$. In addition, $\|I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|=1$.
Next, we present three inequalities in Lemmas 5.14, 5.15, and 5.16, which will assist in deriving our main results. The proofs of these lemmas are provided in Appendix 9.
Lemma 5.14.
When $\alpha_{k+1}\leq\frac{2}{\mu+L}$ for all $k\geq 0$, the following inequality holds for every $k\geq 0$:
$\quad\mathbb{E}\left[\|\bar{\bm{x}}(k+1)-\bm{x}^{*}\|^{2}\right]$ (19)
$\leq(1-\alpha_{k+1}\mu)\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}\right]$
$\quad+\frac{L^{2}\alpha_{k+1}}{n\mu}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}\right]$
$\quad+2p(1+2L\alpha_{k+1}+\frac{L^{2}\alpha_{k+1}}{\mu})\nu_{k+1}^{2}.$
Lemma 5.15.
The following inequality holds for all $k\geq 1$:
$\quad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{1}\bar{\bm{x}}(k+1)\|^{2}\right]$ (20)
$\leq\left(\frac{2+\sigma^{2}}{3}+\frac{4(2+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\right)\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}\right]$
$\quad+\frac{4n(2+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}\right]$
$\quad+\frac{(2+\sigma^{2})(q_{1}^{2}+2L^{2}\alpha_{k+1}^{2})}{1-\sigma^{2}}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}\right]$
$\quad+2np\left(9+6L\alpha_{k+1}+\frac{18L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\right)\nu_{k+1}^{2}$
$\quad+2np\Big{(}15+6L\alpha_{k}+4L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1}$
$\qquad+\frac{6L^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\nu_{k}^{2},$
where $\sigma\triangleq\max_{k\geq 1}\{\|\tilde{W}(k)-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|\}<1$.
Lemma 5.16.
The following inequality holds for all $k\geq 1$:
$\quad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{x}(k)\|^{2}\right]$ (21)
$\leq\Big{(}\frac{3+\sigma^{2}}{4}+\frac{3(3+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}\right]$
$\quad+\frac{4n(3+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}\right]$
$\quad+\frac{(3+\sigma^{2})(\|W-I\|^{2}+4L^{2}\alpha_{k+1}^{2})}{1-\sigma^{2}}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}\right]$
$\quad+2np\Big{(}9+6L\alpha_{k+1}+\frac{28L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\nu_{k+1}^{2}$
$\quad+4np\Big{(}19+6L\alpha_{k}+5L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1}$
$\qquad+\frac{6L^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\nu_{k}^{2}.$
where $\sigma$ is defined as in Lemma 5.15.
Let $s_{1}(k)\triangleq\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}\right]$, $s_{2}(k)\triangleq\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}\right]$, and $s_{3}(k)\triangleq\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}\right]$. In light of Lemmas 5.14, 5.15, and 5.16, we have
$\bm{s}(k+1)\leq A_{k}\bm{s}(k)+\bm{b}_{k},$ (22)
where $\bm{s}(k)$, $A_{k}$, and $\bm{b}_{k}$ are defined in (23) and (24).
$\bm{s}(k)\triangleq\left[\begin{matrix}s_{1}(k)\\ s_{2}(k)\\ s_{3}(k)\end{matrix}\right],\ A_{k}\triangleq\left[\begin{matrix}1-\alpha_{k+1}\mu&\frac{L^{2}\alpha_{k+1}}{n\mu}&0\\ \frac{4n(2+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}&\frac{2+\sigma^{2}}{3}+\frac{4(2+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}&\frac{(2+\sigma^{2})(q_{1}^{2}+2L^{2}\alpha_{k+1}^{2})}{1-\sigma^{2}}\\ \frac{4n(3+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}&\frac{(3+\sigma^{2})(\|W-I\|^{2}+4L^{2}\alpha_{k+1}^{2})}{1-\sigma^{2}}&\frac{3+\sigma^{2}}{4}+\frac{3(3+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\end{matrix}\right],$ (23)
$\bm{b}_{k}\triangleq\left[\begin{matrix}2p(1+2L\alpha_{k+1}+\frac{L^{2}\alpha_{k+1}}{\mu})\nu_{k+1}^{2}\\ 2np\big{(}9+6L\alpha_{k+1}+\frac{18L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\big{)}\nu_{k+1}^{2}+2np\big{(}15+6L\alpha_{k}+4L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1}+\frac{6L^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}}{1-\sigma^{2}}\big{)}\nu_{k}^{2}\\ 2np\big{(}9+6L\alpha_{k+1}+\frac{28L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\big{)}\nu_{k+1}^{2}+4np\big{(}19+6L\alpha_{k}+5L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1}+\frac{6L^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}}{1-\sigma^{2}}\big{)}\nu_{k}^{2}\end{matrix}\right].$ (24)
We now present the convergence result and optimization accuracy of our proposed algorithm in the following theorem.
Theorem 5.17.
Consider the distributed optimization problem (1) with Assumptions 1, 2, 3. If the stepsize decay parameter $q_{1}$ in (5) satisfy $q_{1}^{2}\leq\frac{(1-\sigma^{2})^{4}}{48(\theta+1)(2+\sigma^{2})(3+\sigma^{2})\|W-I\|^{2}}$, where $\sigma<1$ is defined in Lemma 5.15 and $\theta>1$ is a constant, then the sequence $\{\mathbf{x}(k)\}$ generated by Algorithm 1 converges, i.e., $\lim_{k\to\infty}\mathbb{E}\left[\left\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\right\|^{2}\right]=0$. The optimization accuracy $d\triangleq\lim_{k\to\infty}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}\right]$ is bounded by
$e^{-\frac{\mu\gamma}{1-q_{1}}}c_{1}+\frac{\gamma c_{2}L^{2}}{n\mu(1-q_{1})}+\frac{2p\delta^{2}\gamma^{2}q_{2}^{2}}{\epsilon^{2}(q_{2}-q_{1})^{2}(1-q_{2}^{2})}$ (25)
$\quad+\Big{(}4pL+\frac{2pL^{2}}{\mu}\Big{)}\frac{\delta^{2}q_{2}^{2}\gamma^{3}}{\epsilon^{2}(q_{2}-q_{1})^{2}(1-q_{1}q_{2}^{2})},$
where $c_{1},c_{2}>0$ are constants.
Proof 5.18.
We first prove that under appropriate stepsize parameters, $\rho(A_{k})\leq 1$ for any $k\geq 1$. Then, we prove the boundedness of $\bm{s}(k)$ and finally obtain the desired results. Considering that $\alpha_{k}\to 0$ as $k\to\infty$, we analyze the spectral radius of $A_{k}$ both for $k<\infty$ and $k\to\infty$, respectively.
For the case of $k<\infty$, we will demonstrate that $\rho(A_{k})$ is strictly less than 1. By Lemma 5.12, it suffices to ensure that $[A_{k}]_{11},[A_{k}]_{22},[A_{k}]_{33}<1$ and $\det(I-A_{k})>0$. If $\alpha_{k+1}\leq\frac{2}{\mu+L}$, then $10\leq[A_{k}]_{11}=1-\alpha_{k+1}\mu<$ 1. To guarantee that $[A_{k}]_{22}<1$ and $[A_{k}]_{33}<1$, we only need $\alpha_{k+1}^{2}<\frac{(1-\sigma^{2})^{2}}{12(3+\sigma^{2})L^{2}}$. If $\alpha_{k+1}$ further satisfies
$\alpha_{k+1}^{2}$ $\leq\frac{(1-\sigma^{2})^{2}}{24(3+\sigma^{2})L^{2}},$ (26)
then $1-[A_{k}]_{22}\geq\frac{1-\sigma^{2}}{6}$ and $1-[A_{k}]_{33}\geq\frac{1-\sigma^{2}}{8}$. Suppose $\alpha_{k+1}$ also satisfies the following relations:
$\quad[A_{k}]_{12}[A_{k}]_{21}\leq\frac{1}{\theta+1}\left(1-[A_{k}]_{11}\right)\left(1-[A_{k}]_{22}\right),$ (27a)
$\quad[A_{k}]_{23}[A_{k}]_{32}\leq\frac{1}{\theta+1}\left(1-[A_{k}]_{22}\right)\left(1-[A_{k}]_{33}\right),$ (27b)
$\quad[A_{k}]_{12}[A_{k}]_{23}[A_{k}]_{31}$
$\leq\frac{1}{\theta+2}\big{(}\left(1-[A_{k}]_{11}\right)\left(1-[A_{k}]_{22}\right)\left(1-[A_{k}]_{33}\right)$ (27c)
$\quad-[A_{k}]_{12}[A_{k}]_{21}(1-[A_{k}]_{33})-[A_{k}]_{23}[A_{k}]_{32}(1-[A_{k}]_{11})\big{)},$
where $\theta>1$. We then have
$\quad\det(I-A_{k})$ (28)
$=(1-[A_{k}]_{11})(1-[A_{k}]_{22})(1-[A_{k}]_{33})$
$\quad-[A_{k}]_{12}[A_{k}]_{21}(1-[A_{k}]_{33})-[A_{k}]_{23}[A_{k}]_{32}(1-[A_{k}]_{11})$
$\quad-[A_{k}]_{12}[A_{k}]_{23}[A_{k}]_{31}$
$\geq\frac{\theta+1}{\theta+2}\big{(}(1-[A_{k}]_{11})(1-[A_{k}]_{22})(1-[A_{k}]_{33})$
$\qquad-[A_{k}]_{12}[A_{k}]_{21}(1-[A_{k}]_{33})-[A_{k}]_{23}[A_{k}]_{32}(1-[A_{k}]_{11})\big{)}$
$\geq\frac{\theta-1}{\theta+2}(1-[A_{k}]_{11})(1-[A_{k}]_{22})(1-[A_{k}]_{33})$
$>0.$
Therefore, we obtain $\rho(A_{k})<1$ for any $1\leq k<\infty$.
For the case of $k\to\infty$,
$A_{\infty}=\left[\begin{matrix}1&0\\ 0&\tilde{A}\end{matrix}\right],\tilde{A}=\left[\begin{matrix}\frac{2+\sigma^{2}}{3}&\frac{(2+\sigma^{2})q_{1}^{2}}{1-\sigma^{2}}\\ \frac{(3+\sigma^{2})\|W-I\|^{2}}{1-\sigma^{2}}&\frac{3+\sigma^{2}}{4}\end{matrix}\right].$ (29)
If $q_{1}$ satisfies
$q_{1}^{2}<\frac{(1-\sigma^{2})^{4}}{12(2+\sigma^{2})(3+\sigma^{2})\|W-I\|^{2}},$ (30)
then $\det(I-\tilde{A})=(1-\tilde{A}_{11})(1-\tilde{A}_{22})-\tilde{A}_{12}\tilde{A}_{21}>0$. In light of Lemma 5.12, we have $\rho(\tilde{A})<1$.
We now show that $\bm{s}(k)$ is bounded for all $k\geq 1$. By Lemma 5.6.10 in [20], for any $\varepsilon>0$, there exists an induced matrix norm $\|\cdot\|_{s}$ such that
$\|A_{k}\|_{s}\leq\left\{\begin{aligned} &\rho(A_{k})+\varepsilon,\ \ k<\infty,\\ &\max\{1,\rho(\tilde{A})+\varepsilon\},\ \ k\to\infty.\end{aligned}\right.$ (31)
Since $\rho(A_{k})<1$ for $k<\infty$ and $\rho(\tilde{A})<1$, we can choose $\varepsilon$ small enough such that $\|A_{k}\|_{s}\leq 1$ for all $k\geq 1$. Then, by recursively invoking (22), it follows that
$\|\bm{s}(k+1)\|_{s}$ $\leq\|A_{k}A_{k-1}\cdots A_{1}\bm{s}(1)\|_{s}+\sum_{i=1}^{k}\Big{\|}\Big{(}\prod_{j=i+1}^{k}A_{j}\Big{)}\bm{b}_{i}\Big{\|}_{s}$ (32)
$\leq\|\bm{s}(1)\|_{s}+\sum_{i=1}^{\infty}\|\bm{b}_{i}\|_{s}.$
Since $\sum_{i=1}^{\infty}[\bm{b}_{i}]_{j}<\infty$ for $j=1,2,3$ and by the equivalence of norms, we have $\sum_{i=1}^{\infty}\|\bm{b}_{i}\|_{s}<\infty$. Thus, $\bm{s}(k)$ is bounded for all $k\geq 1$.
Let $s_{1}(k)\leq c_{1}$, $s_{2}(k)\leq c_{2}$, and $s_{3}(k)\leq c_{3}$, where $c_{1},c_{2},c_{3}>0$ are some constants. Substituting these bounds into (22), we obtain
$s_{1}(k+1)$ $\leq(1-\alpha_{k+1}\mu)s_{1}(k)+c_{2}[A_{k}]_{12}+[\bm{b}_{k}]_{1},$ (33a)
$\tilde{\bm{s}}(k+1)$ $\leq\tilde{A}\tilde{\bm{s}}(k)+\tilde{\bm{b}}_{k},$ (33b)
where $\tilde{\bm{s}}(k)\triangleq\left[s_{2}(k),s_{3}(k)\right]^{T}$, $\tilde{A}$ is defined as in (29), and $[\tilde{\bm{b}}_{k}]_{1}\triangleq\frac{2(2+\sigma^{2})(2nc_{1}+2c_{2}+c_{3})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}+[\bm{b}_{k}]_{2}$, $[\tilde{\bm{b}}_{k}]_{2}\triangleq\frac{(3+\sigma^{2})(4nc_{1}+4c_{2}+3c_{3})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}+[\bm{b}_{k}]_{3}$.
Since $\rho(\tilde{A})<1$, it follows that $\tilde{\bm{s}}(k)$ converges and satisfies $\tilde{\bm{s}}(\infty)\leq(I-\tilde{A})^{-1}\tilde{\bm{b}}_{\infty}$. Given that $\tilde{\bm{b}}_{\infty}=0$, it follows that $\lim_{k\to\infty}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}\right]=0$. For the optimization accuracy, by (33a) and the parameter settings in (5), we obtain
$\quad\lim_{k\to\infty}\mathbb{E}\left[\left\|\bar{\bm{x}}(k+1)-\bm{x}^{*}\right\|^{2}\right]$ (34)
$\leq\prod_{i=0}^{\infty}(1-\alpha_{i+1}\mu)s_{1}(0)+\sum_{i=0}^{\infty}\prod_{j=i+1}^{\infty}(1-\alpha_{j+1}\mu)\Big{(}\frac{c_{2}L^{2}\alpha_{i+1}}{n\mu}$
$\qquad+2p(1+2L\alpha_{i+1}+\frac{L^{2}\alpha_{i+1}}{\mu})\nu_{i+1}^{2}\Big{)}$
$\leq e^{-\mu\sum_{i=1}^{\infty}\alpha_{i}}c_{1}+\frac{c_{2}L^{2}}{n\mu}\sum_{i=1}^{\infty}\alpha_{i}+2p\sum_{i=1}^{\infty}\nu_{i}^{2}$
$\quad+\Big{(}4pL+\frac{2pL^{2}}{\mu}\Big{)}\sum_{i=1}^{\infty}\alpha_{i}\nu_{i}^{2}$
$\leq e^{-\frac{\mu\gamma}{1-q_{1}}}c_{1}+\frac{\gamma c_{2}L^{2}}{n\mu(1-q_{1})}+\frac{2p\delta^{2}\gamma^{2}q_{2}^{2}}{\epsilon^{2}(q_{2}-q_{1})^{2}(1-q_{2}^{2})}$
$\quad+\Big{(}4pL+\frac{2pL^{2}}{\mu}\Big{)}\frac{\delta^{2}q_{2}^{2}\gamma^{3}}{\epsilon^{2}(q_{2}-q_{1})^{2}(1-q_{1}q_{2}^{2})},$
where the second inequality holds due to $1-x\leq e^{-x}$ for all $x\in\mathbb{R}$.
It remains to show that (27) holds. From (27a), we require
$\alpha_{k+1}^{2}\leq\frac{\mu^{2}(1-\sigma^{2})^{2}}{24(\theta+1)(2+\sigma^{2})L^{4}}.$ (35)
From (27b), we need
$\quad\frac{(2+\sigma^{2})(3+\sigma^{2})(q_{1}^{2}+2L^{2}\alpha_{k+1}^{2})(\|W-I\|^{2}+4L^{2}\alpha_{k+1}^{2})}{(1-\sigma^{2})^{2}}$ (36)
$\leq\frac{(1-\sigma^{2})^{2}}{48(\theta+1)},$
equivalently,
$q_{1}^{2}\leq\frac{(1-\sigma^{2})^{4}}{48(\theta+1)(2+\sigma^{2})(3+\sigma^{2})\|W-I\|^{2}},$ (37)
$\alpha_{k+1}^{2}\leq\frac{-c_{4}+\sqrt{c_{4}^{2}+32c_{5}L^{4}}}{16L^{4}},$ (38)
where $c_{4}\triangleq 4L^{2}q_{1}^{2}+2L^{2}\|W-I\|^{2}$ and $c_{5}\triangleq\frac{(1-\sigma^{2})^{4}}{48(\theta+1)(2+\sigma^{2})(3+\sigma^{2})}-q_{1}^{2}\|W-I\|^{2}$. From (27c), we need
$\quad\frac{4(2+\sigma^{2})(3+\sigma^{2})(q_{1}^{2}+2L^{2}\alpha_{k+1}^{2})L^{4}\alpha_{k+1}^{2}}{\mu(1-\sigma^{2})^{2}}$ (39)
$\leq\frac{\mu(\theta-1)(1-\sigma^{2})^{2}}{48(\theta+1)(\theta+2)},$
which means that
$\alpha_{k+1}^{2}\leq\frac{(-1+\sqrt{1+c_{6}})q_{1}^{2}}{4L^{2}},$ (40)
where $c_{6}\triangleq\frac{\mu^{2}(\theta-1)(1-\sigma^{2})^{4}}{24(\theta+1)(\theta+2)(2+\sigma^{2})(3+\sigma^{2})L^{2}q_{1}^{4}}$. In summary, the stepsize $\alpha_{k}$ must satisfy conditions (26), (35), (38), (40), and $\alpha_{k+1}\leq\frac{2}{\mu+L}$. Since $\alpha_{k}$ decays exponentially, there always exists a $K$ such that $\alpha_{k}$ meets all the aforementioned conditions for all $k\geq K$. Let $k=k-K$. Considering (30) and (37), we complete the proof.
In Theorem 5.17, we first establish the convergence of Algorithm 1, demonstrating that all agents eventually reach a consistent solution in the mean square sense. Subsequently, we derive a bound on the optimization accuracy of Algorithm 1. This accuracy bound depends on four parameters: $\gamma,q_{1},q_{2},\epsilon$. For fixed $\gamma,q_{1},q_{2}$, the accuracy bound has the order of $\mathcal{O}(\frac{1}{\epsilon^{2}})$. It reveals the inherent trade-off between exact convergence and DP: the smaller the privacy budget $\epsilon$, the higher the privacy level, but the lower the accuracy.
Remark 5.19.
The established bound (25) shows the trade-off between privacy and accuracy. Superior optimization accuracy can be achieved under a given privacy budget by solving the following optimization problem:
$\min_{\gamma,q_{1},q_{2}}d(\gamma,q_{1},q_{2})$ (41)
subject to $\gamma>0,q_{1}\in(0,1),q_{2}\in(q_{1},1).$
Note that Problem (41) is non-convex, which poses challenges for finding the global optimum. To address this issue, one possible approach is to utilize non-convex optimizers. Alternatively, a heuristic strategy can be employed to obtain a locally optimal solution: 1) Randomly initialize $\gamma$, $q_{1}$, and $q_{2}$ within their feasible domains. 2) Fix two parameters and optimize the remaining one to a local optimum. 3) Repeat Step 2) by iteratively optimizing different parameters several times.
6 Numerical Results
In this section, we demonstrate the effectiveness of Algorithm 1 using the canonical sensor fusion problem. In this problem, all $n$ sensors collaborate to estimate an unknown variable $\bm{x}$ by solving the following optimization problem:
$\min_{\bm{x}\in\mathbb{R}^{p}}\sum_{i=1}^{n}\left(\|\bm{v}_{i}-M_{i}\bm{x}\|^{2}+\omega_{i}\|\bm{x}\|^{2}\right),$ (42)
where $\bm{v}_{i}\in\mathbb{R}^{m}$, $M_{i}\in\mathbb{R}^{m\times p}$, and $\omega_{i}\in\mathbb{R}$ represent the observation, measurement matrix, and regularization coefficient of sensor $i$, respectively. The parameter settings are consistent with those in [10].
Figure 1: The communication topology of 100 sensors.
Figure 2: Comparison of convergence between Algorithm 1 and the DPOP method [12] under the same privacy budget.
Figure 3: Comparison of convergence between Algorithm 1 and the PP-DOAGT method [16] under the same noise.
Table 3: Comparison of the Privacy, Accuracy, Applicability, and Efficiency of the Three Different Types of Algorithms.
Methods M-NMI under M-NMI under Accuracy Topological Efficiency
external eavesdropper honest-but-curious adversary assumption
Algorithm 1 with $\epsilon=10$ 0.52 0.52 $1.9\times 10^{-4}$ High
Algorithm 1 with $\epsilon=1$ 0.24 0.24 $2.0\times 10^{-3}$ High
Algorithm 1 with $\epsilon=0.1$ 0.047 0.047 $3.0\times 10^{-2}$ High
Correlated randomness-based method [6] 1 0.47 $6.6\times 10^{-9}$ High
Encryption-based method [9] 0 0.30 $3.2\times 10^{-6}$ Low
6.1 Comparison with Other Differentially Private Methods
We consider a multi-agent system of 100 sensors, where the communication network is generated using an Erdős-Rényi random graph model with an edge creation probability of 0.1, as illustrated in Fig. 1. The dimensions in Problem (42) are set as $m=3$ and $p=2$. For Algorithm 1, the initial states $\bm{x}_{i}(0)$ are randomly sampled from a standard normal distribution. The stepsize $\alpha_{k}$ and noise parameter $\nu_{k}$ are set according to (5), and the relevant parameters $\gamma$, $\beta$, $q_{2}$, and $q_{1}$ are tuned following the heuristic strategy described in Remark 5.19. We implement Algorithm 1 under different privacy budgets ($\epsilon\in\{0.1,1,10\}$). For $\epsilon=0.1/1/10$, the parameters are set as $\gamma=0.001/0.001/0.002$, $\beta=1000/1000/100$, $q_{2}=0.99/0.99/0.99$, and $q_{1}=0.92/0.97/0.97$, respectively. For each case, we calculate the mean and standard deviation of the residuals (denoted as ${\|\mathbf{x}(k)-\mathbf{x}^{*}\|^{2}}$) over 100 trials, with the results shown as the red line in Fig. 2. It can be observed that our algorithm quickly approaches the optimal solution under all privacy budgets. This confirms that the proposed algorithm can achieve arbitrary $\epsilon$-DP. Moreover, higher privacy levels result in lower optimization accuracy, which aligns with the conclusion of Theorem 5.17.
For comparison, we implement the DPOP method from [12] under the same privacy budgets. The corresponding results are shown by the blue lines in Fig. 2. It can be seen that our algorithm achieves higher accuracy than the DPOP method. This is because Algorithm 1’s sensitivity is more compact, and the introduction of auxiliary variable $\bm{y}_{i}$ enhances its convergence. Additionally, we also compare the proposed algorithm with the latest PP-DOAGT method from [16] under the same noise conditions. The comparison results are shown in Fig. 3. At higher privacy levels (the privacy budgets of the PP-DOAGT method are 18446.44, 3652.67, and 365.27 with $\delta=1$, respectively), our algorithm achieves comparable optimization accuracy to the PP-DOAGT method. This is due to the fact that our algorithm has lower sensitivity than the PP-DOAGT method (as shown in Table 2).
Figure 4: Schematic illustration of the honest-but-curious adversary and the external eavesdropper.
6.2 Comparison with Other Types of Privacy-Preserving Methods
To objectively evaluate the privacy levels of different types of privacy-preserving algorithms, we employ the Maximum Normalized Mutual Information (M-NMI), a modified mutual information-based metric from [21], as the privacy measure:
[…]
where $\mathcal{V}_{i}(k)$ is the private variable of agent $i$ at iteration $k$, $\mathcal{I}_{j}(k)$ represent the information attacker $j$ can obtain about $\mathcal{V}_{i}(k)$, and $K$ is the total number of iterations. $I(\mathcal{V}_{i}(k),\mathcal{I}_{j}(k))$ is the mutual information between $\mathcal{V}_{i}(k)$ and $\mathcal{I}_{j}(k)$, which quantifies the amount of agent $i$’s private information contained in the information acquired by attacker $j$. $I(\mathcal{V}_{i}(k),\mathcal{V}_{i}(k))$ represents the entropy of agent $i$’s private information, which is used for normalization. M-NMI quantifies the worst-case privacy leakage across all iterations. A smaller M-NMI value indicates stronger privacy preservation of the algorithm. For $\mathcal{V}_{i}(k)$, we consider the agent’s gradients as its private information [2]. For $\mathcal{I}_{j}(k)$, it depends on the attacker’s model and the algorithm’s update rule. We consider two common attacker models: the honest-but-curious adversary and the external eavesdropper [22]. We adopt the system as shown in Fig. 4, in which agent 1 is an honest agent, agents 2 and 3 are honest-but-curious adversaries, and agent $e$ is an external eavesdropper. Agents 1, 2, and 3 collaboratively solve Problem (42) with dimensions $m=1$ and $p=1$. Agents 2, 3, and $e$ all attempt to infer the private information of agent 1. For each method, we perform 5,000 Monte Carlo trials, each comprising 1,000 iterations. We then compute its M-NMI values under the honest-but-curious adversary and the external eavesdropper using the NPEET toolbox [23].
Under Algorithm 1, the private information of agent 1 is $\nabla f_{1}\left(\bm{z}_{1}(k)\right),1\leq k\leq K$, i.e., $\mathcal{V}_{1}(k)=\left\{\nabla f_{1}\left(\bm{z}_{1}(k)\right)\right\}$. Honest-but-curious adversaries 2 and 3 can collude to infer the private information of agent 1. They receive $\bm{z}_{1}(k)$ from agent 1, infer $\bm{y}_{1}(k)$ using update (4b) and public information $W$ and $\beta$, and approximate $\nabla f_{1}\left(\bm{z}_{1}(k)\right)$ using update (4c) and public information $\alpha_{k}$, that is,
$\mathcal{I}_{2}(k)=\left\{\bm{z}_{1}(k),\bm{y}_{1}(k),\frac{1}{\alpha_{k}}\left(\bar{\bm{z}}_{1}(k)-\bm{z}_{1}(k)\right)-\bm{y}_{1}(k)\right\}.$ (44)
The external eavesdropper $e$ acquires the same information about agent 1 as the honest-but-curious adversaries, leading to $\mathcal{I}{e}(k)=\mathcal{I}_{2}(k)$. The corresponding simulation results under different privacy budgets are shown in the first three rows of Table 3. It can be seen that our algorithm can simultaneously defend against the honest-but-curious adversary and the external eavesdropper. Furthermore, as the privacy budget decreases, both the M-NMI value and optimization accuracy decline, which aligns with the intuition of a differentially private algorithm.
For comparison, we implement the correlated randomness-based privacy-preserving method from [6]. The correlated randomness-based method relies on an additional topological assumption: each honest agent has at least one neighboring agent that is also honest. Thus, for this method, we assume that agent 3 in Fig. 4 is also an honest agent, and we compute its M-NMI under the honest-but-curious adversary using the following information:
$\mathcal{I}_{2}(k)=\left\{\bm{x}_{1}^{k},\Lambda_{1}^{k}\bm{y}_{1}^{k},C_{21}^{k}\bm{y}_{1}^{k}+B_{21}^{k}(\nabla f_{1}(\bm{x}_{1}^{k+1})-\nabla f_{1}(\bm{x}_{1}^{k}))\right\},$
which consists of information directly received by agent 2 from agent 1. The relevant results are shown in the fourth row of Table 3. It can be observed that this method achieves high accuracy, as the correlated randomness-based method enables exact convergence. Under the external eavesdropper, the M-NMI value of this method is 1, indicating that it cannot defend against the external eavesdropper. This is because the privacy of the correlated randomness-based method relies on the additional topological assumption, which does not work against the external eavesdropper. We also implement the encryption-based privacy-preserving method from [9]. Its M-NMI under the honest-but-curious adversary is computed using the following information:
$\mathcal{I}_{2}(k)=\Big{\{}\frac{1}{\alpha_{k}}\Big{(}\sum_{j\in\{2,3\}}b_{1\to j}b_{j\to 1}(\bm{x}_{j}(k)-\bm{x}_{1}(k)\Big{)}\Big{\}},$
which is an estimate of agent 1’s gradient according to its update rule. The corresponding results are shown in the last row of Table 3. It can be seen that this method achieves an M-NMI value of 0 under the external eavesdropper, meaning the encryption-based method is fully resistant to the external eavesdropper. This is because the external eavesdropper can only access encrypted data. Moreover, this method exhibits high accuracy as the encryption-based method can also achieve exact convergence. However, the encryption-based method requires numerous encryption and decryption operations, resulting in low computational and communication efficiency. In addition, although all three types of algorithms can defend against the honest-but-curious adversary, Table 3 demonstrates that our algorithm allows for enhanced privacy performance through adjustable privacy budgets, whereas the privacy performance of the other two methods remains fixed.
7 Conclusion
In this paper, we propose a differentially private distributed optimization algorithm. Compared to existing state-of-the-art algorithms, the proposed algorithm exhibits lower sensitivity, thereby achieving higher optimization accuracy under the same privacy budget. Furthermore, the proposed algorithm reduces both the number of communication variables and the amount of injected noise by half, significantly enhancing communication and computational efficiency. Moreover, we rigorously analyze the algorithm’s differential privacy, establish its convergence, and derive an upper bound on its optimization accuracy. The theoretical results reveal the inherent trade-off between privacy and accuracy and provide guidance for parameter tuning. Future work will focus on developing differentially private distributed algorithms with lower sensitivity for directed graphs.
[…]
8 Proof of Proposition 1
Proof 8.20.
We prove the proposition by directly comparing the algorithm’s sensitivity under different designs. We consider the classical differentially private distributed gradient descent (DP-DGD) as the baseline algorithm:
$\mathbf{x}(k)=W\mathbf{z}(k)-\alpha_{k}\nabla F(\mathbf{z}(k)),$ (45)
where $\mathbf{z}(k)\triangleq\mathbf{x}(k-1)+\bm{\xi}(k)$ is the noisy state and $\alpha_{k}$ is the stepsize. Note that in the DP-DGD algorithm, agents use their own noisy states for consensus and compute gradients based on noisy states. According to Definition 3, and since
$\|\mathbf{x}(k)-\mathbf{x}^{\prime}(k)\|_{1}$ $=\|\bm{x}_{i_{0}}(k)-\bm{x}_{i_{0}}^{\prime}(k)\|_{1}$ (46)
$=\alpha_{k}\|\nabla f_{i_{0}}(\bm{z}_{i_{0}}(k))-\nabla f_{i_{0}}^{\prime}(\bm{z}_{i_{0}}(k))\|_{1}$
$\leq\delta\alpha_{k},$
the sensitivity of the DP-DGD algorithm, denoted as $\Delta_{1}(k)$, is bounded by $\delta\alpha_{k}$, i.e., $\Delta_{1}(k)\leq\delta\alpha_{k}$. The proof details for (46) can be found in Theorem 4.5 or in Lemma 4 of [12].
If agents use their true states for consensus, the DP-DGD algorithm can be reformulated as
$\mathbf{x}(k)=W_{0}\mathbf{x}(k-1)+(W-W_{0})\mathbf{z}(k)-\alpha_{k}\nabla F(\mathbf{z}(k)),$ (47)
where $W_{0}\triangleq\text{diag}\{W\}$. Let $\Delta_{2}(k)$ denote the sensitivity of the above algorithm at iteration $k$. Since $\|\bm{x}_{i_{0}}(k)-\bm{x}_{i_{0}}^{\prime}(k)\|_{1}=\|W_{i_{0}i_{0}}(\bm{x}_{i_{0}}(k-1)-\bm{x}_{i_{0}}^{\prime}(k-1))-\alpha_{k}(\nabla f_{i_{0}}(\bm{z}_{i_{0}}(k))-\nabla f_{i_{0}}^{\prime}(\bm{z}_{i_{0}}(k)))\|_{1}\leq W_{i_{0}i_{0}}\Delta_{2}(k-1)+\delta\alpha_{k}$, we have $\Delta_{2}(k)\leq W_{i_{0}i_{0}}\Delta_{2}(k-1)+\delta\alpha_{k}$. Due to $\Delta_{2}(k-1)\geq 0$ for all $k\geq 1$, we obtain the first argument.
If agents compute gradients based on their true states, the DP-DGD algorithm is reformulated as
$\mathbf{x}(k)=W\mathbf{z}(k)-\alpha_{k}\nabla F(\mathbf{x}(k-1)).$ (48)
Let $\Delta_{3}(k)$ denote the sensitivity of the above algorithm at iteration $k$. Since $\|\bm{x}_{i_{0}}(k)-\bm{x}_{i_{0}}^{\prime}(k)\|_{1}=\alpha_{k}\|\nabla f_{i_{0}}(\bm{x}_{i_{0}}(k-1))-\nabla f_{i_{0}}^{\prime}(\bm{x}_{i_{0}}^{\prime}(k-1))\|_{1}=\alpha_{k}\|\nabla f_{i_{0}}(\bm{x}_{i_{0}}(k-1))-\nabla f_{i_{0}}(\bm{x}_{i_{0}}^{\prime}(k-1))+\nabla f_{i_{0}}(\bm{x}_{i_{0}}^{\prime}(k-1))-\nabla f_{i_{0}}^{\prime}(\bm{x}_{i_{0}}^{\prime}(k-1))\|_{1}\leq\sqrt{n}L\alpha_{k}\Delta_{3}(k-1)+\delta\alpha_{k}$, we obtain $\Delta_{3}(k)\leq\sqrt{n}L\alpha_{k}\Delta_{3}(k-1)+\delta\alpha_{k}$. Because of $\Delta_{3}(k-1)\geq 0$ for all $k\geq 1$, we have the second argument.
For the third argument, when neighboring agents share only variable $\bm{x}$, the algorithm’s sensitivity depends on the bound of $\|\mathbf{x}(k)-\mathbf{x}^{\prime}(k)\|_{1}$. If agents are also required to exchange additional variables (e.g., an auxiliary variable $\bm{y}$), the algorithm’s sensitivity increases to $\|\mathbf{x}(k)-\mathbf{x}^{\prime}(k)\|_{1}+\|\mathbf{y}(k)-\mathbf{y}^{\prime}(k)\|_{1}$. Hence, the third argument holds.
9 Proof of Basic Lemmas
9.1 Proof of Lemma 5.10
Proof 9.21.
It follows from Assumption 3 that $\mathbb{E}\left[\bm{\xi}(k)|\mathcal{M}_{k}\right]=0$, $\mathbb{E}\left[\|\bm{\xi}(k)\||\mathcal{M}_{k}\right]\leq np\nu_{k}$, and $\mathbb{E}\left[\|\bm{\xi}(k)\|^{2}|\mathcal{M}_{k}\right]=2np\nu_{k}^{2}$. For part (a), it follows from the definition of $\mathbf{z}(k)$ that
$\mathbb{E}\left[\langle\mathbf{z}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$ $=\mathbb{E}\left[\langle\bm{\xi}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$ (49)
$\leq\mathbb{E}\left[\|\bm{\xi}(k)\|^{2}|\mathcal{M}_{k}\right]$
$=2np\nu_{k}^{2}.$
For part (b), in light of Assumption 1, we have
$\quad\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (50)
$=\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k))-\nabla F(\mathbf{x}(k-1)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\leq L\mathbb{E}\left[\|\bm{\xi}(k)\|^{2}|\mathcal{M}_{k}\right]$
$=2npL\nu_{k}^{2}.$
For part (c), following a similar derivation as in Lemma 8 of [19], we obtain
$\quad\mathbb{E}\left[\left\langle\nabla F(\mathbf{x}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (51)
$\leq\sum_{i=1}^{n}\left(L|W_{ii}-\alpha_{k}\beta+\alpha_{k}\beta W_{ii}|+L^{2}\alpha_{k}\right)\mathbb{E}\left[\|\bm{\xi}_{i}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\leq 2np(L+L^{2}\alpha_{k})\nu_{k}^{2},$
where the second inequality follows from the fact that $W_{ii}-\alpha_{k}\beta+\alpha_{k}\beta W_{ii}\in(-1,1)$. Then, we have
$\quad\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k+1)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (52)
$=\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k+1))-\nabla F(\mathbf{x}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[\left\langle\nabla F(\mathbf{x}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\leq L\mathbb{E}\left[\|\bm{\xi}(k+1)\|\|\bm{\xi}(k)\||\mathcal{M}_{k}\right]+2np(L+L^{2}\alpha_{k})\nu_{k}^{2}$
$\leq n^{2}p^{2}L\nu_{k+1}\nu_{k}+2np(L+L^{2}\alpha_{k})\nu_{k}^{2}$
$\leq npL(2+np+2L\alpha_{k})\nu_{k}^{2}.$
For part (d), by substituting (13a) into (13b), we get
$\mathbf{x}(k)$ $=((1+\alpha_{k}\beta)W-\alpha_{k}\beta I)\mathbf{z}(k)-\alpha_{k}\mathbf{y}(k-1)$ (53)
$\quad-\alpha_{k}\nabla F(\mathbf{z}(k)).$
Since $\mathbf{y}(k-1)$ is uniquely determined by $\mathcal{M}_{k}$, and based on parts (a) and (b), we have
$\quad\mathbb{E}\left[\langle\mathbf{x}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$ (54)
$=\mathbb{E}\left[\left\langle((1+\alpha_{k}\beta)W-\alpha_{k}\beta I)\mathbf{z}(k)-\alpha_{k}\nabla F(\mathbf{z}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\leq 2np\left(\|(1+\alpha_{k}\beta)W-\alpha_{k}\beta I\|+L\alpha_{k}\right)\nu_{k}^{2}$
$\leq 2np\left(3+L\alpha_{k}\right)\nu_{k}^{2}.$
For part (e), it follows from (14) and parts (a) and (b) that
$\quad\mathbb{E}\left[\langle\mathbf{1}\bar{\bm{x}}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$ (55)
$=\mathbb{E}\left[\left\langle\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\mathbf{z}(k)-\frac{\alpha_{k}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\leq 2np(1+L\alpha_{k})\nu_{k}^{2}.$
By the definition of $\mathbf{z}(k)$, part (f) can be obtained as follows:
$\quad\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$ (56)
$=\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)+\bm{\xi}(k+1)\|^{2}|\mathcal{M}_{k}\right]$
$=\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{\xi}(k+1)\|^{2}|\mathcal{M}_{k}\right]$
$=\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]+2np\nu_{k+1}^{2}.$
Similarly, we have
$\quad\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{z}(k)\|^{2}|\mathcal{M}_{k}\right]$ (57)
$=\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)+\bm{\xi}(k+1)-\bm{\xi}(k)\|^{2}|\mathcal{M}_{k}\right]$
$=\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{\xi}(k+1)-\bm{\xi}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2\mathbb{E}\left[\langle\mathbf{x}(k),-\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$
$\leq\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]+2np\nu_{k+1}^{2}$
$\quad+2np(7+2L\alpha_{k})\nu_{k}^{2},$
where in the first inequality we use part (d).
9.2 Proof of Lemma 5.14
Proof 9.22.
Let $\bm{a}_{1}\triangleq\bar{\bm{x}}(k)-\frac{\alpha_{k+1}}{n}\mathbf{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))-\bm{x}^{*}$, $\bm{b}_{1}\triangleq\frac{\alpha_{k+1}}{n}\mathbf{1}^{T}\left(\nabla F(\mathbf{z}(k+1))-\nabla F(\mathbf{1}\bar{\bm{x}}(k))\right)$, and $\bm{c}_{1}\triangleq\frac{1}{n}\mathbf{1}^{T}\bm{\xi}(k+1)$. In light of (15) and Lemma 5.9, if $\alpha_{k+1}\leq\frac{2}{\mu+L}$ for any $k\geq 0$, we have
$\quad\mathbb{E}\left[\|\bar{\bm{x}}(k+1)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$ (58)
$=\mathbb{E}\left[\|\bm{a}_{1}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{b}_{1}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{1}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[2\langle\bm{a}_{1},-\bm{b}_{1}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{b}_{1},\bm{c}_{1}\rangle|\mathcal{M}_{k}\right]$
$\leq(1-\alpha_{k+1}\mu)^{2}\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{L^{2}\alpha_{k+1}^{2}}{n}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2p\nu_{k+1}^{2}+\mathbb{E}\left[2\langle\bm{a}_{1},-\bm{b}_{1}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{b}_{1},\bm{c}_{1}\rangle|\mathcal{M}_{k}\right].$
We now need to bound the last two terms in the above inequality. By Lemma 5.9, we obtain
$\quad 2\langle\bm{a}_{1},-\bm{b}_{1}\rangle$ (59)
$\leq 2\Big{\|}\bar{\bm{x}}(k)-\frac{\alpha_{k+1}}{n}\mathbf{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))-\bm{x}^{*}\Big{\|}\|\bm{b}_{1}\|$
$\leq 2(1-\alpha_{k+1}\mu)\|\bar{\bm{x}}(k)-\bm{x}^{*}\|\|\bm{b}_{1}\|$
$\leq(1-\alpha_{k+1}\mu)\Big{(}\alpha_{k+1}\mu\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}+\frac{1}{\alpha_{k+1}\mu}\|\bm{b}_{1}\|^{2}\Big{)}$
$\leq\alpha_{k+1}\mu(1-\alpha_{k+1}\mu)\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}$
$\quad+\frac{L^{2}\alpha_{k+1}(1-\alpha_{k+1}\mu)}{n\mu}\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2},$
where in the third inequality we use Young’s inequality. Then, taking the conditional expectation, we get
$\quad\mathbb{E}\left[2\langle\bm{a}_{1},-\bm{b}_{1}\rangle|\mathcal{M}_{k}\right]$ (60)
$\leq\alpha_{k+1}\mu(1-\alpha_{k+1}\mu)\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{L^{2}\alpha_{k+1}(1-\alpha_{k+1}\mu)}{n\mu}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right].$
For the last term in (58), since $\mathbf{x}(k)$ is independent of $\bm{\xi}(k+1)$, we have
$\quad\mathbb{E}\left[2\langle-\bm{b}_{1},\bm{c}_{1}\rangle|\mathcal{M}_{k}\right]$ (61)
$=\frac{2\alpha_{k+1}}{n^{2}}\mathbb{E}\left[\left\langle-\mathbf{1}^{T}\nabla F(\mathbf{z}(k+1)),\mathbf{1}^{T}\bm{\xi}(k+1)\right\rangle|\mathcal{M}_{k}\right]$
$\leq 4pL\alpha_{k+1}\nu_{k+1}^{2},$
where the first inequality follows from the part (b) of Lemma 5.10. Finally, combining (58), (60), and (61) gives
$\quad\mathbb{E}\left[\|\bar{\bm{x}}(k+1)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$ (62)
$\leq(1-\alpha_{k+1}\mu)\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{L^{2}\alpha_{k+1}}{n\mu}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+(2p+4pL\alpha_{k+1})\nu_{k+1}^{2}$
$\leq(1-\alpha_{k+1}\mu)\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{L^{2}\alpha_{k+1}}{n\mu}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2p(1+2L\alpha_{k+1}+\frac{L^{2}\alpha_{k+1}}{\mu})\nu_{k+1}^{2},$
where the second inequality follows from the part (f) of Lemma 5.10. Taking full expectation on both sides of the above inequality completes the proof.
9.3 Proof of Lemma 5.15
Proof 9.23.
Let $\bm{a}_{2}\triangleq\tilde{W}(k)\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)$, $\bm{b}_{2}\triangleq\frac{\alpha_{k+1}}{\alpha_{k}}W(\mathbf{x}(k)-\mathbf{x}(k-1))$, $\bm{c}_{2}\triangleq\alpha_{k+1}(\nabla F(\mathbf{z}(k+1))-\nabla F(\mathbf{z}(k)))-\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k+1))$, and $\bm{d}_{2}\triangleq\left((1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\right)\bm{\xi}(k+1)-\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)$. In light of (17), we have
$\quad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{1}\bar{\bm{x}}(k+1)\|^{2}|\mathcal{M}_{k}\right]$ (63)
$=\mathbb{E}\left[\|\bm{a}_{2}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{b}_{2}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{2}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[\|\bm{d}_{2}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{2},\bm{b}_{2}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{2},-\bm{c}_{2}\rangle|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[2\langle\bm{b}_{2},-\bm{c}_{2}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[2\langle\bm{b}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{c}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right].$
Next, we will bound each term on the right-hand side of the above equality, respectively. From Remark 5.8, it follows that $\tilde{W}(k)$ is doubly-stochastic and $\sigma_{k}=\|\tilde{W}(k)-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|<1$ for all $k\geq 1$. Thus, we have
$\mathbb{E}\left[\|\bm{a}_{2}\|^{2}|\mathcal{M}_{k}\right]$ $=\mathbb{E}\Big{[}\Big{\|}\Big{(}\tilde{W}(k)-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\Big{)}(\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$ (64)
$\leq\sigma^{2}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right],$
where $\sigma\triangleq\max_{k\geq 1}\{\sigma_{k}\}<1$. For the second term in (63), we have
$\mathbb{E}\left[\|\bm{b}_{2}\|^{2}|\mathcal{M}_{k}\right]$ $\leq q_{1}^{2}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right].$ (65)
It follows from Lemma 5.9 that
$\quad\Big{\|}\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{z}(k+1))\Big{\|}^{2}$ (66)
$=\Big{\|}\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{z}(k+1))-\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))+\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))\Big{\|}^{2}$
$\leq 2\Big{\|}\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{z}(k+1))-\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))\Big{\|}^{2}$
$\quad+2\Big{\|}\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))\Big{\|}^{2}$
$\leq\frac{2L^{2}}{n}\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}+2L^{2}\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}.$
Then, we have
$\quad\mathbb{E}\left[\|\bm{c}_{2}\|^{2}|\mathcal{M}_{k}\right]$ (67)
$\leq 2\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\nabla F(\mathbf{z}(k+1))-\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2n\alpha_{k+1}^{2}\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{z}(k+1))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\leq 2L^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\mathbf{z}(k+1)-\mathbf{z}(k)\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+4L^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+4nL^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}|\mathcal{M}_{k}\right]$
$\leq 2L^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+4L^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+4nL^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+12npL^{2}\alpha_{k+1}^{2}\nu_{k+1}^{2}+4npL^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}\nu_{k}^{2},$
where the last inequality follows from the parts (f) and (g) of Lemma 5.10.
For the fourth term of (63), since $\bm{\xi}(k+1)$ and $\bm{\xi}(k)$ are independent of each other, we have
$\mathbb{E}\left[\|\bm{d}_{2}\|^{2}|\mathcal{M}_{k}\right]$ $\leq 9\mathbb{E}\left[\left\|\bm{\xi}(k+1)\right\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\left\|\bm{\xi}(k)\right\|^{2}|\mathcal{M}_{k}\right]$ (68)
$=18np\nu_{k+1}^{2}+2np\nu_{k}^{2}.$
Similar to the proof in (64), it follows that
$\quad\mathbb{E}\left[2\langle\bm{a}_{2},\bm{b}_{2}\rangle|\mathcal{M}_{k}\right]$ (69)
$\leq 2\sigma\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|\|\bm{b}_{2}\||\mathcal{M}_{k}\right]$
$\leq\sigma\Big{(}\frac{1-\sigma^{2}}{3\sigma}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]+\frac{3\sigma}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{b}_{2}\|^{2}|\mathcal{M}_{k}\right]\Big{)}$
$=\frac{1-\sigma^{2}}{3}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]+\frac{3\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{b}_{2}\|^{2}|\mathcal{M}_{k}\right],$
where in the second inequality we use Young’s Inequality.
Similarly, we have
$\quad\mathbb{E}\left[2\langle\bm{a}_{2},-\bm{c}_{2}\rangle|\mathcal{M}_{k}\right]$ (70)
$\leq\frac{1-\sigma^{2}}{3}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]+\frac{3\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{c}_{2}\|^{2}|\mathcal{M}_{k}\right],$
and
$\mathbb{E}\left[2\langle\bm{b}_{2},-\bm{c}_{2}\rangle|\mathcal{M}_{k}\right]$ $\leq\mathbb{E}\left[\|\bm{b}_{2}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{2}\|^{2}|\mathcal{M}_{k}\right].$ (71)
For the last three terms of (63), it follows from Lemma 5.10 that
$\mathbb{E}\left[2\langle\bm{a}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]$ $=\frac{2\alpha_{k+1}}{\alpha_{k}}\mathbb{E}\left[\Big{\langle}\tilde{W}(k)\mathbf{x}(k),-W\bm{\xi}(k)\Big{\rangle}|\mathcal{M}_{k}\right]$ (72)
$\quad+\frac{2\alpha_{k+1}}{\alpha_{k}}\mathbb{E}\left[\left\langle\mathbf{1}\bar{\bm{x}}(k),W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\leq 8np(2+L\alpha_{k})\nu_{k}^{2},$
$\mathbb{E}\left[2\langle\bm{b}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]$ $=\frac{2\alpha_{k+1}^{2}}{\alpha_{k}^{2}}\mathbb{E}\left[\left\langle W\mathbf{x}(k),-W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (73)
$\leq 4np(3+L\alpha_{k})\nu_{k}^{2},$
$\quad\mathbb{E}\left[2\langle-\bm{c}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]$ (74)
$=2\alpha_{k+1}\mathbb{E}\Big{[}\Big{\langle}(\frac{1}{n}\mathbf{1}\mathbf{1}^{T}-I)\nabla F(\mathbf{z}(k+1)),$
$\qquad\Big{(}(1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\Big{)}\bm{\xi}(k+1)\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\quad+2\frac{\alpha_{k+1}^{2}}{\alpha_{k}}\mathbb{E}\Big{[}\Big{\langle}(\frac{1}{n}\mathbf{1}\mathbf{1}^{T}-I)\nabla F(\mathbf{z}(k+1)),-W\bm{\xi}(k)\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\quad+2\frac{\alpha_{k+1}^{2}}{\alpha_{k}}\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k)),-W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\leq 12npL\alpha_{k+1}\nu_{k+1}^{2}+2npL(4+np+2L\alpha_{k})\alpha_{k+1}\nu_{k}^{2}.$
Finally, combining (63)-(65) and (67)-(74), we have
$\quad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{1}\bar{\bm{x}}(k+1)\|^{2}|\mathcal{M}_{k}\right]$ (75)
$\leq\frac{2+\sigma^{2}}{3}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]+\frac{2+\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{b}_{2}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{2+\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{c}_{2}\|^{2}|\mathcal{M}_{k}\right]+2np(9+6L\alpha_{k+1})\nu_{k+1}^{2}$
$\quad+2np(15+6L\alpha_{k}+4L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1})\nu_{k}^{2}$
$\leq\Big{(}\frac{2+\sigma^{2}}{3}+\frac{4(2+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{4n(2+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{(2+\sigma^{2})(q_{1}^{2}+2L^{2}\alpha_{k+1}^{2})}{1-\sigma^{2}}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2np\Big{(}9+6L\alpha_{k+1}+\frac{18L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\nu_{k+1}^{2}$
$\quad+2np\Big{(}15+6L\alpha_{k}+4L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1}$
$\qquad+\frac{6L^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\nu_{k}^{2}.$
Taking the full expectation for the above inequality gives the desired result.
9.4 Proof of Lemma 5.16
Proof 9.24.
In light of (18), and let $\bm{a}_{3}\triangleq\frac{\alpha_{k+1}}{\alpha_{k}}W(\mathbf{x}(k)-\mathbf{x}(k-1))$, $\bm{b}_{3}\triangleq(1-\frac{\alpha_{k+1}}{\alpha_{k}}+\alpha_{k+1}\beta)\left(W-I\right)\mathbf{x}(k)$, $\bm{c}_{3}\triangleq\alpha_{k+1}(\nabla F(\mathbf{z}(k+1))-\nabla F(\mathbf{z}(k)))$, $\bm{d}_{3}\triangleq\left((1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I\right)\bm{\xi}(k+1)-\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)$, we have
$\quad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{x}(k)\|^{2}|\mathcal{M}_{k}\right]$ (76)
$=\mathbb{E}\left[\|\bm{a}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[\|\bm{d}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{3},\bm{b}_{3}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{3},-\bm{c}_{3}\rangle|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[2\langle\bm{b}_{3},-\bm{c}_{3}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[2\langle\bm{b}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{c}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right].$
Similar to the proof of Lemma 5.15, we will bound each term on the right-hand side of the above equality. For the first term, it follows from (14) that
$\quad W(\mathbf{x}(k)-\mathbf{x}(k-1))$ (77)
$=\Big{(}W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\Big{)}(\mathbf{x}(k)-\mathbf{x}(k-1))+\mathbf{1}\bar{\bm{x}}(k)-\mathbf{1}\bar{\bm{x}}(k-1)$
$=\Big{(}W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\Big{)}(\mathbf{x}(k)-\mathbf{x}(k-1))-\frac{\alpha_{k}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))$
$\quad+\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k).$
Then, we obtain
$\quad\mathbb{E}\left[\|\bm{a}_{3}\|^{2}|\mathcal{M}_{k}\right]$ (78)
$\leq\sigma^{2}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k)\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\quad+\alpha_{k+1}^{2}\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\quad+2q_{1}\mathbb{E}\Big{[}\Big{\langle}(W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T})(\mathbf{x}(k)-\mathbf{x}(k-1)),$
$\qquad-\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\quad+2q_{1}^{2}\mathbb{E}\Big{[}\Big{\langle}(W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T})\mathbf{x}(k),\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k)\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\quad+2q_{1}\mathbb{E}\Big{[}\Big{\langle}-\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k)),\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k)\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\leq\sigma^{2}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\Big{[}\Big{\|}\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\quad+\sigma\Big{(}\frac{1-\sigma^{2}}{4\sigma}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+\frac{4\sigma}{1-\sigma^{2}}\mathbb{E}\Big{[}\left\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\Big{]}\Big{)}$
$\quad+2np(7+2L\alpha_{k}+2L\alpha_{k+1})\nu_{k}^{2}$
$=\frac{1+3\sigma^{2}}{4}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{1+3\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\Big{[}\Big{\|}\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\quad+2np(7+2L\alpha_{k}+2L\alpha_{k+1})\nu_{k}^{2},$
where the first inequality is due to the fact that $\|W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|\leq\sigma<1$, and the second inequality follows from Young’s Inequality as well as parts (b) and (d) of Lemma 5.10.
For the second term in (76), we have
$\quad\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]$ (79)
$=\left(1-q_{1}+\alpha_{k+1}\beta\right)^{2}\mathbb{E}\left[\left\|(W-I)(\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k))\right\|^{2}|\mathcal{M}_{k}\right]$
$\leq\|W-I\|^{2}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right].$
By part (g) of Lemma 5.10, we obtain
$\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right]$ $\leq L^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{z}(k)\|^{2}|\mathcal{M}_{k}\right]$ (80)
$\leq L^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2npL^{2}\alpha_{k+1}^{2}\nu_{k+1}^{2}+2npL^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}\nu_{k}^{2}.$
For the fourth term in (76), since $\bm{\xi}(k+1)$ and $\bm{\xi}(k)$ are independent of each other, we have
$\mathbb{E}\left[\|\bm{d}_{3}\|^{2}|\mathcal{M}_{k}\right]$ $=\mathbb{E}\left[\left\|\left((1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I\right)\bm{\xi}(k+1)\right\|^{2}|\mathcal{M}_{k}\right]$ (81)
$\quad+\frac{\alpha_{k+1}^{2}}{\alpha_{k}^{2}}\mathbb{E}\left[\left\|W\bm{\xi}(k)\right\|^{2}|\mathcal{M}_{k}\right]$
$\leq 18np\nu_{k+1}^{2}+2np\nu_{k}^{2}.$
For the fifth term in (76), it follows from (77) that
$\quad\mathbb{E}\left[2\langle\bm{a}_{3},\bm{b}_{3}\rangle|\mathcal{M}_{k}\right]$ (82)
$=2q_{1}\mathbb{E}\Big{[}\Big{\langle}(W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T})(\mathbf{x}(k)-\mathbf{x}(k-1)),\bm{b}_{3}\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\quad+2\mathbb{E}\left[\left\langle-\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k)),\bm{b}_{3}\right\rangle|\mathcal{M}_{k}\right]$
$\quad+2q_{1}\mathbb{E}\Big{[}\Big{\langle}\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k),(\tilde{W}(k)-I)\mathbf{x}(k)\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\leq\sigma\Big{(}\frac{1-\sigma^{2}}{4\sigma}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]+\frac{4\sigma}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]\Big{)}$
$\quad+\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\Big{[}\Big{\|}\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\quad+8np(3+L\alpha_{k})\nu_{k}^{2}$
$=\frac{1-\sigma^{2}}{4}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]+\frac{1+3\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\Big{[}\Big{\|}\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}+8np(3+L\alpha_{k})\nu_{k}^{2},$
where the first inequality follows from Young’s Inequality and part (d) of Lemma 5.10.
Similar to the proof in (82), we have
$\quad\mathbb{E}\left[2\langle\bm{a}_{3},-\bm{c}_{3}\rangle|\mathcal{M}_{k}\right]$ (83)
$=2q_{1}\mathbb{E}\Big{[}\Big{\langle}(W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T})(\mathbf{x}(k)-\mathbf{x}(k-1)),-\bm{c}_{3}\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\quad+2\mathbb{E}\left[\left\langle\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k)),\bm{c}_{3}\right\rangle|\mathcal{M}_{k}\right]$
$\quad+2q_{1}\alpha_{k+1}\mathbb{E}\Big{[}\Big{\langle}\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k),\nabla F(\mathbf{z}(k))-\nabla F(\mathbf{z}(k+1))\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\leq\sigma\Big{(}\frac{1-\sigma^{2}}{4\sigma}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]+\frac{4\sigma}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right]\Big{)}$
$\quad+\mathbb{E}\left[\left\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2npL(4+np+2L\alpha_{k})\alpha_{k+1}\nu_{k}^{2}$
$=\frac{1-\sigma^{2}}{4}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]+\frac{1+3\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[\left\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2npL(4+np+2L\alpha_{k})\alpha_{k+1}\nu_{k}^{2},$
where the first inequality follows from the parts (b) and (c) of Lemma 5.10. In addition,
$\mathbb{E}\left[2\langle\bm{b}_{3},-\bm{c}_{3}\rangle|\mathcal{M}_{k}\right]\leq\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right].$ (84)
For the last three terms in (76), we use Lemma 5.10 to obtain
$\mathbb{E}\left[2\langle\bm{a}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]$ $=2\mathbb{E}\left[\left\langle\frac{\alpha_{k+1}}{\alpha_{k}}W\mathbf{x}(k),-\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (85)
$\leq 4np(3+L\alpha_{k})\nu_{k}^{2},$
$\quad\mathbb{E}\left[2\langle\bm{b}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]$ (86)
$=2\mathbb{E}\left[\left\langle(\tilde{W}(k)-I)\mathbf{x}(k),-\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\leq 8np(3+L\alpha_{k})\nu_{k}^{2},$
$\quad\mathbb{E}\left[2\langle-\bm{c}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]$ (87)
$=2\mathbb{E}[\langle-\alpha_{k+1}\nabla F(\mathbf{z}(k+1)),$
$\qquad\left((1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I\right)\bm{\xi}(k+1)\rangle|\mathcal{M}_{k}]$
$\quad+2\mathbb{E}\left[\left\langle\alpha_{k+1}\nabla F(\mathbf{z}(k+1)),\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\quad+2\mathbb{E}\left[\left\langle\alpha_{k+1}\nabla F(\mathbf{z}(k)),-\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\leq 12npL\alpha_{k+1}\nu_{k+1}^{2}+2npL(4+np+2L\alpha_{k})\alpha_{k+1}\nu_{k}^{2}.$
To derive the final result, we also need to bound $\mathbb{E}\left[\left\|\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\right]$. From (66), we have
$\quad\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$ (88)
$\leq 2\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))-\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k+1))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\quad+2\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k+1))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\leq\frac{2L^{2}}{n}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{z}(k)\|^{2}|\mathcal{M}_{k}\right]+4L^{2}\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{4L^{2}}{n}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\leq\frac{2L^{2}}{n}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]+4L^{2}\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{4L^{2}}{n}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]+12pL^{2}\nu_{k+1}^{2}$
$\quad+4pL^{2}(7+2L\alpha_{k})\nu_{k}^{2},$
where the third inequality follows from parts (f) and (g) of Lemma 5.10.
Finally, by combining (76) and (78)-(88), we obtain
$\quad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{x}(k)\|^{2}|\mathcal{M}_{k}\right]$ (89)
$\leq\frac{3+\sigma^{2}}{4}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]+6np(3+2L\alpha_{k+1})\nu_{k+1}^{2}$
$\quad+4np(19+6L\alpha_{k}+5L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1})\nu_{k}^{2}$
$\quad+\frac{3+\sigma^{2}}{1-\sigma^{2}}\Big{(}\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+n\alpha_{k+1}^{2}\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}\Big{)}$
$\leq\Big{(}\frac{3+\sigma^{2}}{4}+\frac{3(3+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{4n(3+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{(3+\sigma^{2})(\|W-I\|^{2}+4L^{2}\alpha_{k+1}^{2})}{1-\sigma^{2}}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2np\Big{(}9+6L\alpha_{k+1}+\frac{28L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\nu_{k+1}^{2}$
$\quad+4np\Big{(}19+6L\alpha_{k}+5L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1}$
$\qquad+\frac{6L^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\nu_{k}^{2}.$
Taking the full expectation on both sides of the above inequality completes the proof.